### v1

# FIG: Forward-Inverse Generation for Low-Resource Domain-specific Event Detection

Abstract
Event Detection (ED) is the task of identifying typed event mentions of interest from natural language text, which benefits domain-specific reasoning in biomedical, legal, and epidemiological domains. However, procuring supervised data for thousands of events for various domains is a laborious and expensive task. To this end, existing works have explored synthetic data generation via forward (generating labels for unlabeled sentences) and inverse (generating sentences from generated labels) generations. However, forward generation often produces noisy labels, while inverse generation struggles with domain drift and incomplete event annotations. To address these challenges, we introduce FIG, a hybrid approach that leverages inverse generation for high-quality data synthesis while anchoring it to domain-specific cues extracted via forward generation on unlabeled target data. FIG further enhances its synthetic data by adding missing annotations through forward generation-based refinement. Experimentation on three ED datasets from diverse domains reveals that FIG outperforms the best baseline, achieving average gains of 3.3% F1 and 5.4% F1 in the zero-shot and few-shot settings, respectively. Analyzing the generated trigger hit rate and human evaluation substantiates FIG’s superior domain alignment and data quality compared to existing baselines. We will release our code at https://github.com/PlusLabNLP/FIG.
FIG: Forward-Inverse Generation for Low-Resource
Domain-specific Event Detection
Tanmay Parekh      Yuxuan Dong      Lucas Bandarkar      Artin Kim I-Hung Hsu      Kai-Wei Chang      Nanyun Peng Computer Science Department, University of California, Los Angeles        Google {tparekh, violetpeng, kwchang}@cs.ucla.edu
1 Introduction
Figure 1: Comparing different data generation paradigms. Owing to poor LLM ED reasoning skills, forward generation (top) struggles with noisy labels. Inverse generation (middle) suffers from domain drift owing to lack of domain-specific cues. We introduce FIG (bottom), a hybrid forward and inverse approach, to overcome both issues to generate better quality data.
Event Detection (ED) Sundheim (1992); Doddington et al. (2004) involves identifying and categorizing significant events from natural language text based on a pre-defined ontology. It has applications in widespread domains like biomedicine Pyysalo et al. (2012), epidemiology Parekh et al. (2024b, c), law Francesconi et al. (2010), etc. However, procuring annotations for each domain-specific ontology to train models is expensive and impractical. Although recent works introduce zero-shot LLM-reasoning Gao et al. (2023); Cai et al. (2024), their performance falls short of supervised approaches Huang et al. (2024).
To mitigate the need for extensive annotations to train supervised models, prior work has explored LLM-powered synthetic data generation to enhance downstream model training. A widely used approach is a forward generation or weak supervision He et al. (2021); Chia et al. (2022), where labels are assigned to existing unlabeled data (i.e., $X\rightarrow Y^{\prime}$). This generation is limited by LLM ED reasoning, often leading to noisy labels. To address this, inverse generation has been proposed Meng et al. (2022); Wang et al. (2023c), where sentences are generated based on sampled or generated labels (i.e $Y\rightarrow X^{\prime}$). However, inverse generation introduces a domain drift between synthetic and target data owing to the lack of any domain-specific cues, as illustrated in Figure 1.
In our work, we propose Forward-Inverse Generation (FIG), a hybrid approach that integrates forward and inverse generation to enhance synthetic data quality. Instead of generating event triggers from scratch — an inherently high-variance process — we first apply forward generation to unlabeled data, extracting domain-specific triggers. Next, we utilize inverse generation to synthesize diverse sentences conditioned on the domain-specific triggers. Finally, a second forward generation annotates any missing events in the generated data to ensure high data quality. This approach produces cleaner synthetic data compared to forward generation while maintaining a closer domain alignment with target data than inverse generation, as shown in Figure 1.
We evaluate FIG on ED datasets from three different domains: ACE Doddington et al. (2004) (news), SPEED Parekh et al. (2024c) (social media), and GENIA2011 Kim et al. (2011) (biomedical). Our primary method of evaluation is testing ED performance of DEGREE Hsu et al. (2022) trained on the generated data. In the zero-shot setting, FIG outperforms STAR Ma et al. (2024) (inverse generation baseline) and weak supervision (forward generation baseline) by an average of 16.8% and 3.5% Tri-C F1, respectively. Similarly, in the few-shot setting, FIG surpasses the strongest baseline by an average of 6.2% Tri-C F1. We study the generated trigger hit rate (relative to the gold trigger set) and demonstrate how FIG’s triggers have about 7.3% better hit rate compared to the best baseline, which contributes to its superior performance. Finally, we also conduct human evaluation, which supports FIG’s superior domain relevance and data annotation quality.
2 Problem Definition
We focus on the task of Event Detection (ED) Sundheim (1992); Doddington et al. (2004) for this work. The task of ED aims to extract mentions of any events of interest from natural language text. Following ACE 2005 Doddington et al. (2004), we define an event as something that happens or describes a change of state and is labeled by a specific event type. The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is referred to as the event mention. Event Detection, in particular, requires extracting the event triggers from the sentence and classifying it into one of the pre-defined event types. We provide an illustration of this task below where arrested and campaigns are the trigger words for the event types Justice: Arrest-Jail and Conflict: Demonstrate respectively.
Some 3,000 people have been arrested since the disobedience campaigns began last week.
[…]
In our work, we specifically focus on ED in diverse and specialized domains (e.g., biomedical), where procuring a training dataset $D_{T}$ of annotated data points is expensive, whereas reasonable-sized unlabeled data $D_{T}^{\prime}$ is available. We focus on two realistic low-resource data setups - zero-shot (zero labeled data) and few-shot ($k$ labeled datapoints per event type) settings. Different from domain transfer, we do not consider any labeled data for the source domain and directly optimize model performance for the target domain.
Figure 2: Model Architecture Diagram. Top left - illustration of forward generation. Bottom left - illustration of inverse generation. Right - FIG and its four components. We first extract domain-specific triggers via forward generation, then generate passages using inverse generation. Forward generation refines missing events, and we sample $N$ data points per event for downstream training.
3 Methodology
In this work, we focus on Large Language Model (LLM)-powered synthetic data generation to alleviate the need for domain-specific annotated training data, as zero-shot LLM-based approaches usually struggle in ED in specialized domains Huang et al. (2024). By generating a large amount of data instances $D_{s}=\{(X,Y)\}$, we can train downstream supervised ED models. Here, we first briefly review prior works on forward generation and backward generation. Next, we introduce our proposed data generation method FIG.
3.1 Background
We focus on two major paradigms for data generation in our work: forward generation and inverse generation. We describe them briefly below and provide illustrations in Figure 2.
Forward Generation:
This is a more straightforward manner of data generation, wherein LLMs are utilized to generate labels $Y^{\prime}$ on unlabeled data $X\in D_{T}^{\prime}$ (i.e. $X\rightarrow Y^{\prime})$. This generation is analogous to weak/distant supervision (Weak Sup) Mintz et al. (2009); Wang et al. (2021); He et al. (2021) where noisy/weak labels are assigned to clean sentences and then utilized to train downstream models. Consequently, the label quality is dependent on the reasoning capability of the LLM.
Inverse Generation:
Inverse generation is a relatively new data generation paradigm Kumar et al. (2020); Schick and Schütze (2021). This paradigm first generates/samples potential labels $Y$ based on the task definition and then generates plausible $X^{\prime}$ conditioned on the label $Y$ (i.e. $Y\rightarrow X^{\prime}$). Recent works like SynthIE Josifoski et al. (2023) and STAR Ma et al. (2024) have explored the usage of LLM-guided inverse generation for information extraction and event extraction tasks. Inverse generation provides control over data distribution via prompting, but the data quality is dependent on the LLM’s sentence generation capability.
3.2 FIG
LLM reasoning has been poor for ED Li et al. (2023a); Huang et al. (2024), which makes forward generation vulnerable to noisy label quality. Since LLMs are pre-trained to generate natural sentences, their generation capabilities are stronger, which supports inverse generation. However, owing to the lack of any domain-specific information, inverse generation can synthesize highly diverse sentences, leading to a domain drift. Merely specifying the domain in the prompt is not sufficient (shown in § D.1). Furthermore, inverse generation sentences can mention additional events which remain unannotated introducing noise in the data.
To this end, we propose Forward-Inverse Generation (FIG), a hybrid approach that infuses forward generation reasoning into the base pipeline of inverse generation to ensure closer alignment to the target domain while improving the quality of the annotated labels. To procure domain-specific cues, we assume access to an unlabeled target domain data $D_{T}^{\prime}$ (similar to forward generation). We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
Target Structure Extraction:
Similar to inverse generation, we first create potential triggers (i.e. labels) for each event type. However, generation from scratch can lead to very divergent set of triggers. For example, an “Attack" event can refer to a war in the news, cyberattacks in the cybersecurity domain, diseases in epidemiology, or criticism in economics. Naturally, each scenario would assume different triggers, which inverse generation struggles to generate even when provided with clear event definitions (actual examples in § 6.5).
Instead, we generate potential triggers per event type by utilizing an LLM to extract triggers using a forward generation over the unlabeled data $D_{T}^{\prime}$. To ensure highly precise extraction of triggers, we develop a two-stage prompt setup. The first stage is tasked to identify and filter possible event types mentioned in the text based on the task instructions and event definitions. The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type. We illustrate this setup in § A.
After extraction, we aggregate and sort the triggers for each event type at the corpus level and filter out the top $t$ triggers for each event type as the set of clean triggers. This filtering helps remove many noisy triggers introduced in the forward generation. Finally, we create target structures $Y$ by sampling 1-2 event types per instance and sampling triggers from the clean trigger set for each sampled event type. We ensure uniform sampling of event types and triggers to avoid unbalanced label distribution. Overall, such forward generation-based trigger extraction ensures the distillation of domain-specific knowledge in the target structures.
Passage Generation:
This component is the core of inverse generation and is tasked with generating passages corresponding to the constraints of the sampled target structure. Specifically, the LLM is prompted with the task instructions, the event definitions, the sampled target structure $Y$, and asked to generate a passage $X^{\prime}$ which mentions the event types using the triggers in $Y$. We illustrate this prompt in Figure 8. The key to generating better target data aligned passages lies in the presence of the target data aligned target structure $Y$ (shown via examples in § 6.5).
Data Refinement and Sampling:
While passage generation ensures the target structure events are mentioned in the passage, the passage could still have other unknown events which can lead to under-annotated $(X^{\prime},Y)$ data instances. We provide an illustration in Figure 3 with the target trigger positive, but the generated sentence also has a missing mention of symptom event triggered by got. To account for such missing annotations, we utilize forward generation to identify all possible event mentions in the generated passage. We remove duplicates and add the new event mentions to create a more complete target structure $Y_{f}$.
Missed Event Trigger: “got” Event: symptom Generation Trigger: “Positive” Event: infect Ok, I just got a fever… Theres a possibility Im COVID-19 Positive
Figure 3: Illustration of how inverse generation can produce unannotated event mentions. Blue box = target event mention, red box = unannotated event mention.
To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger. Additionally, we standardize trigger annotations by correcting variations in trigger word forms. Finally, we apply a greedy sampling algorithm to sample $N$ instances $(X^{\prime},Y_{f})$ per event type to create our final synthetic dataset $D_{s}$.
Downstream Model Training:
The final component utilizes the generated synthetic data $D_{s}$ to train downstream ED models in a supervised manner. The trained ED models are then used for inferring on the test set and for eventual evaluation.
4 Experimental Setup
Base LLM Method Unlabeled ACE SPEED GENIA Average
Data Source Eve-I Tri-C Eve-I Tri-C Eve-I Tri-C Eve-I Tri-C
Llama3-8B Inference - 30.2 23.8 39.8 25.4 21.9 17.2 30.6 22.1
STAR - 44.9 35.0 21.0 10.1 25.9 19.0 30.6 21.4
Weak Sup train 41.7 37.8 45.6 31.5 26.9 21.4 38.1 30.2
FIG(ours) train 57.4 50.2 44.6 31.5 35.2 28.9 45.7 36.9
FIG (ours) external 57.7 52.6 47.8 32.9 33.6 24.6 46.4 36.7
Llama3-70B Inference - 46.9 41.3 46.9 35.6 34.2 28.2 42.7 35.0
STAR - 50.0 42.3 18.3 13.8 23.3 16.9 30.5 24.3
Weak Sup train 53.2 48.0 52.8 39.6 36.2 29.1 47.4 38.9
FIG (ours) train 58.1 53.8 49.9 38.7 38.0 29.7 48.7 40.7
FIG (ours) external 59.7 55.6 50.1 39.2 39.2 31.5 49.7 42.1
GPT-3.5 Inference - 33.0 26.2 44.2 32.9 31.2 24.7 36.1 27.9
STAR - 45.0 36.6 21.3 14.6 21.8 14.3 29.4 21.8
Weak Sup train 49.7 44.6 50.7 37.5 37.7 30.1 46.1 37.4
FIG (ours) train 54.8 48.3 50.3 36.8 39.3 31.1 48.1 38.7
FIG (ours) external 54.0 48.5 50.1 36.1 38.7 29.4 47.6 38.0
- Gold Data - 64.6 61.6 64.0 53.5 51.3 44.0 60.0 53.0
Table 1: Zero-shot results comparing FIG with other baselines across three datasets and three base LLMs. Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al. (2022) model trained on data generated by each technique. Eve-I: Event Identification F1, Tri-C: Trigger Classification F1.
Datasets:
We consider three ED datasets from diverse domains for our experiments: (1) ACE Doddington et al. (2004), in the news domain, (2) SPEED Parekh et al. (2024c), in the social media domain, and (3) GENIA Kim et al. (2011), in the biomedical domain. We simplify GENIA by converting the original document level annotations to sentence-level annotations. For the few-shot setting, we compile $k$ data instances from the training data as the few-shot examplers.
For our unlabeled data, we consider two sources: (1) Train - annotation-free training splits (i.e. only the text) of each dataset and (2) External - unlabeled data from other external sources. For ACE, we utilize News Category Dataset Misra (2022) comprising Huffpost news articles from 2012-2022 as the external data source. We filter articles corresponding to political, financial, and business articles. For SPEED, we utilize COVIDKB Zong et al. (2022) mining tweets from the Twitter COVID-19 Endpoint released in April 2020 as the external data source. Finally, we utilize GENIA2013 dataset Kim et al. (2013) as the external data for GENIA. We provide statistics about these datasets in Table 8.
Baseline methods:
We consider three LLM-based techniques for low-resource ED as the baselines for our work. (1) Inference Gao et al. (2023): LLMs are used to directly infer on the target test data using their reasoning capability. (2) STAR Ma et al. (2024): This model is the state-of-the-art inverse generation model for ED. It utilizes trigger generation, passage generation, and data refinement steps without using any unlabeled data, (3) Weak Supervision (Weak Sup): This model acts as the forward generation baseline. We utilize the Inference model to provide labels for the unlabeled data. For an upper bound reference, we also include a human generation baseline (Human) wherein we sample $N$ data instances from the gold training data of each dataset to train the downstream ED model.
Base models:
For our base LLMs, we consider three instruction-tuned LLMs of varying sizes, namely Llama3-8B-Instruct (8B model), Llama3-70B-Instruct (70B model) Dubey et al. (2024), and GPT-3.5 (175B model) Brown et al. (2020). For our downstream ED model, we consider a specialized low-resource model DEGREE Hsu et al. (2022), a generative model prompted to fill event templates powered on a BART-large pre-trained language model Lewis et al. (2020).
Evaluation:
Our primary evaluation metric is the synthesized data-trained model’s ED performance on the final test splits of each dataset. We consider two low-resource settings - zero-shot (no labeled data) and few-shot ($k$ datapoints per event type are used). Note this is different cross-dataset works Cai et al. (2024) which train and test on an exclusive set of event types. For Inference, the LLM is directly run on the test set to procure model predictions. We report the F1 scores for two metrics Ahn (2006) for measuring model performance: (1) Event Identification (Eve-I) - correct identification of events, and (2) Trigger Classification (Tri-C) - correct identification of trigger-event pairs.
Implementation Details:
We follow STAR for the implementation of the baseline models and the majority of hyperparameter settings. For FIG’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation. We generate $N=50$ datapoints per event type for each generation strategy. All our experimental results are reported over an average of three runs. Additional implementation details are provided in Appendix C.
Figure 4: Few-shot results comparing FIG with other baselines across three datasets using Llama3-8B-Instruct as the base LLM. Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al. (2022) model trained on data generated by each technique. Tri-C: Trigger Classification F1, #: Number of.
5 Results
We present the results and insights for our zero-shot and few-shot settings below.
5.1 Zero-shot Results
We present the main results comparing the various methods across different LLMs and datasets in Table 1. For FIG, we show results using both the train data as well as the external data as the external data source. We highlight some of our main findings below.
Forward generation > Inverse generation:
Although STAR performs decently for ACE, its performances for SPEED and GENIA are quite poor (even below the Inference baseline). Since SPEED (social media) and GENIA (biomedical) are more domain-specific than ACE (news), we attribute this poor performance to STAR’s inability to generate domain-specific data instances (further validated in § 6.3). On the other hand, Weak Sup outperforms STAR, providing average gains of 13% Tri-C F1, thus establishing the superiority of forward generation-based synthetic data.
FIG performs the best:
Our hybrid approach combines the benefits from both forward and inverse generation and provides the best performance. On average, FIG beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points - proving how domain-specific cues from unlabeled data can help generate better-aligned inverse generated data. Compared to forward generation Weak Sup baseline, FIG provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1 - suggesting how data diversity and cleaner label quality can help improve model performance.
External data source is effective:
Assuming access to the training data as the unlabeled data source can be a strong assumption and bias for FIG. To cross-validate this assumption’s impact, we also consider FIG with external data sources. Surprisingly, as seen from Table 1, FIG with external data provides similar gains of 4% Eve-I F1 and 3.4% Tri-C F1 over the best baseline. This demonstrates that instead of getting biased by external data, FIG utilizes the useful domain-specific signals (in the form of extracted triggers) from the external data and provides similar/better gains as with using the training data.
Generation ability doesn’t scale up as reasoning ability:
Although FIG performs better across all LLMs, the gains are much higher for Llama3-8B model relative to the Llama3-70B or GPT3.5 models. Comparing the improvements in model performance when scaling up from Llama3-8B to LLama3-70B, inverse generation-centric methods (STAR, FIG) improve by an average of 2.5% F1 points while reasoning-centric methods (Inference, Weak Sup) improve considerably better by an average of 10.8% F1 points. A similar disparity in performance improvements is observed for GPT3.5 as well. This sheds light on how the generation capabilities of LLMs do not scale compared to reasoning capabilities. Simultaneously, it shows that FIG is particularly effective when used with smaller LLMs whose the reasoning capability are poor.
5.2 Few-shot Results
We also study the various methods in the presence of small annotated data as part of our few-shot experiments. Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized. Majorly, we utilize the $k$ shots as in-context examples in the LLM prompts where applicable and append these few-shot examples to the synthesized training data as well. Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$ shot examples. We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 4 and summarize our major findings below.
FIG consistently performs the best:
Similar to zero-shot results, we observe that FIG consistently beats all other baseline models. On an average, FIG outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
Generation abilities improve significantly:
Contrary to our findings of small improvements when scaling up LLM model size in § 5, we observe that inverse generation-centric methods (STAR, FIG) consistently improve as we increase the number of shots. Contrastingly, reasoning-centric models stagnate and don’t improve as much. This highlights how inverse generation becomes particularly effective in the presence of few exemplars.
6 Analysis
In this section, we provide various analyses to study FIG’s superior performance. Unless specified, we utilize Llama3-8B-Instruct as the base LLM for the analyses.
Method ACE SPEED GENIA
FIG 50.2 31.5 28.9
   – Trigger Generation 43.2 27.8 28.2
   – Data Refinement 47.4 23.3 22.0
Table 2: Ablation study for FIG’s various components measured as Tri-C F1 performance across datasets.
Method ACE SPEED GENIA
EI TC EI TC EI TC
FIG 57.4 50.2 44.6 31.5 35.2 28.9
Weak Sup + STAR 46.9 38.9 44.5 29.5 30.2 24.3
Table 3: Comparing FIG’s hybrid approach with data-mixing of forward and inverse generation.
Method ACE SPEED GENIA
STAR 9.6% 15.3% 15.1%
Weak Sup 19.1% 38.4% 44.8%
FIG 23.2% 49.4% 52.5%
Table 4: Reporting the hit rate of synthetic data triggers relative to gold test triggers for the three methods.
6.1 Ablation study
Table 2 shows the ablation study for each of our forward generation components. We observe how forward generation is critical in both trigger generation and data refinement stages with average performance reductions of 3.8% F1 and 6% F1 on removing the respective components from FIG.
6.2 Comparison with Data mixing
Data-mixing Hoffmann et al. (2022); Xie et al. (2023) is a widely used technique for leveraging complementary information across datasets to promote robust downstream model training. We mix forward (Weak Sup) and inverse generation (STAR) as a hybrid baseline to compare with FIG. To keep comparisons fair, we consider $N/2$ data instances per event type from each dataset. Results from Table 3 demonstrate how FIG beats the data-mixing based hybrid model by 5-6% F1, highlighting the need for explicit model design to combine the benefits of forward and inverse generation.
6.3 Analyzing trigger quality and drift
ED models have a strong tendency to over-rely on lexical relations between triggers and events Tong et al. (2022); Ma et al. (2024). Thus, we compare the synthetic data triggers with the gold test triggers as a raw study on the quality and drift of triggers in the synthesized data. Specifically, we extract triggers per event type in both datasets and measure the hit rate of the synthesized triggers on the gold set, as reported in Table 4. As observed, the poor hit rate shows the poor overlap of STAR’s triggers with the gold triggers, which is a primary reason for its domain drift. Furthermore, the consistently stronger precision of FIG relative to Weak Sup demonstrate FIG’s better trigger annotation quality.
Method Naturalness Event Annotation
Relevance Quality
STAR 3.1 3.4 3.1
Weak Sup 4.2 - 2.9
FIG 3.6 4.0 3.6
Table 5: Human evaluation for sentence naturalness, relevance of event in generated sentence, and the annotation quality. 1 = worst, 5 = best.
Dataset Event Method Trigger Sentence
ACE Attack STAR raid As the rebels embarked on a daring trek across the desert, they launched a surprise raid on the heavily guarded fortress, catching the enemy off guard.
FIG shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
SPEED Death STAR asphyxiation The hiker’s life was tragically cut short as asphyxiation occurred after she became stuck in the narrow cave crevice.
FIG killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
GENIA Binding STAR merge The regulatory protein’s ability to activate a specific region of the DNA triggers the merge of two proteins, leading to the modification of gene expression.
FIG bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
Table 6: Qualitative examples demonstrating STAR and FIG’s trigger and sentence generation quality.
Method ACE SPEED GENIA
EI TC EI TC EI TC
FIG 57.4 50.2 44.6 31.5 35.2 28.9
   + SFT LLM 55.2 51.7 46.9 35.8 36.7 29.1
Table 7: Measuring model performance improvement using a LLM fine-tuned on unlabeled train data (SFT LLM) for FIG. EI: Event Identification F1, TC: Trigger Classification F1.
To further study the label quality and relevance, we conduct a human evaluation. Specifically, a human expert ED annotator is tasked to score the generations (between 1-5) on the naturalness of the sentence for the domain, the correct relevance of the event mentioned in the generated sentence, and the annotation quality (more details in § D.3). We provide the averaged scores across the three datasets for 90 samples in Table 5. Weak Sup has high sentence quality but poor label annotations; while STAR suffers from poor event relevance indicating domain drift. Overall, FIG performs the best with high annotation quality and event relevance.
6.4 Domain-adapted LLM Fine-tuning
We fine-tune the base LLM used for passage generation on the unlabeled target-domain data $D_{T}^{\prime}$ to better align the generated passages to the target domain. Naturally, this can be applied only for smaller LLMs owing to fine-tuning costs. We present the results of fine-tuning Llama3-8B-Instruct on the unlabeled train data in Table 7. On average, we observe how target data fine-tuning improves FIG by a slight 0.5-2% F1, suggesting that target-domain passage generation may help, but it is not a highly influencing factor to improve downstream model performance.
6.5 Qualitative analysis of generated data
We provide qualitative evidence for FIG’s domain-aligned triggers compared to STAR in Table 6 (more examples in Table 16). Owing to lack of domain grounding in STAR, the resulting triggers often appear misaligned (e.g. asphyxiation for death event related to pandemics). This misalignment carries over to the generated sentences, further reducing their quality and alignment. In contrast, FIG’s triggers are better aligned to the target domain corpus, resulting in better quality data.
7 Related Works
Low-resource Event Detection
Event Detection (ED) has been studied extensively Sundheim (1992); Grishman and Sundheim (1996), leading to diverse datasets in news Doddington et al. (2004); Song et al. (2015); Ellis et al. (2015), Wikipedia Li et al. (2021); Pouran Ben Veyseh et al. (2022), and general domains Wang et al. (2020); Parekh et al. (2023), as well as niche areas like biomedical Pyysalo et al. (2012); Kim et al. (2011, 2013), multimedia Li et al. (2020), cybersecurity Satyapanich et al. (2020), epidemiology Parekh et al. (2024b, c), and pharmacovigilance Sun et al. (2022). To address the growing need for event detection across expanding domains, better low-resource domain-specific techniques are essential. Prior works have explored transfer learning via Abstract Meaning Representation Huang et al. (2018), Semantic Role Labeling Zhang et al. (2021), and Question Answering Lyu et al. (2021). Reformulating ED as a conditional generation task has also aided low-resource training Hsu et al. (2022, 2023); Huang et al. (2022). Recently, LLM-based reasoning Li et al. (2023a); Gao et al. (2023); Wang et al. (2023b) and transfer-learning Cai et al. (2024) has been explored for low-resource ED and transfer learning, but performance remains inferior to supervised models Huang et al. (2024). This motivates efforts in synthetic data generation for low-resource ED.
Data Generation for Information Extraction
LLM-powered synthetic data generation has been successful for various NLP tasks Li et al. (2023b); Wang et al. (2023c); Wu et al. (2024); Shao et al. (2025). For information extraction, works have explored knowledge retrieval Chen and Feng (2023); Amalvy et al. (2023), translation Parekh et al. (2024a); Le et al. (2024), and data re-editing Lee et al. (2021); Hu et al. (2023). Some works in the directions we focus in our work include forward generation Chia et al. (2022); Ye et al. (2022); Wang et al. (2023a); Tang et al. (2023) and inverse generation Josifoski et al. (2023); Ma et al. (2024). We improve on existing forward and inverse generation works via our hybrid generation approach FIG.
8 Conclusion and Future Work
We introduce FIG, a hybrid forward-inverse generation approach for better domain-specific synthetic data in low-resource ED. Experiments on three diverse datasets reveal that forward generation suffers from noisy labels due to poor LLM reasoning, while inverse generation faces domain drift. FIG mitigates these issues, achieving superior domain alignment and cleaner data, leading to the best downstream ED performance in zero and few-shot settings. Future works can explore enhancing stronger domain alignment and extending FIG for multilingual generation.
Acknowledgments
We express our gratitude to Anh Mac, Po-Nien Kung, and Christina Chance for their valuable time, reviews of our work, and constructive feedback.
Limitations
We consider only Event Detection (ED) as the main task for data generation, but our method can be extended to other structured prediction tasks as well. We leave this exploration for future works. We consider three specialized domains of news, social media, and biomedical to provide a proof-of-concept of our work. There are other specialized domains for ED as well which can be explored as part of future work. Finally, our proposed method FIG makes a practical assumption of access to unlabeled data to procure target domain cues to guide the data generation. However, for specific super-specialized domains or if data has privacy concerns, this may not be possible and our method may not be applicable here. We assume such cases to be super rare and beyond the scope of our work.
Ethical Considerations
The theme of our work is to generate high-quality domain-specific data using Large Language Models (LLMs) using forward-inverse generations. The inherent LLMs can have certain biases, which can lead to potentially harmful or biased generations. Furthermore, the LLM can introduce potential hallucinations in the annotations which can hurt the model performance. We do not check or consider any bias/hallucination detection method as part of our work, as it is beyond the scope. So future users should take due consideration of this vulnerability.
Our proposed method FIG utilizes unlabeled data as a basis to procure domain-specific cues. If there are any biases in this data, it can propogate to the downstream model as well. We provide a proof-of-concept about our method in this work but do not detect or rectify such biases.
Since the inverse generation paradigm causes the LLM to generate sentences/passages, which can potentially be copied from the pre-training data the LLM has been trained on. This can potentially lead to copyright infringements and we do not consider any of such violations under consideration for our method. Users should consider this vulnerability before usage in commercial applications.
We would also like to mention and acknowledge that we have utilized AI chatbots to help with the writing of the work.
Appendix A Additional details about FIG
We provide illustrations about the various prompts used in FIG. For target structure extraction forward generation, we consider a two-prompt approach. The first prompt aims to identify if any events of interest are mentioned in the text as illustrated in Figure 6. It comprises the task definition, full event ontology with definitions, the task instructions, and the query sentence. The second prompt identifies the trigger corresponding to the event of interest, as illustrated in Figure 7. Here, we specify the task definition, the event ontology details, and the query text with the task instructions. To aid inverse generation for passage generation, we provide the task definition, the event ontology with event definitions, and the query comprising the sampled target structure. We illustrate this prompt in Figure 8. Finally, we provide the simplified one-prompt setup for forward generation utilized for data refinement in Figure 9.
Appendix B Additional Experimental Setup Details
B.1 Data Statistics
We discuss details about our dataset in § 4. Our test target domain data includes the test data splits of (1) ACE Doddington et al. (2004) in the news domain, (2) SPEED Parekh et al. (2024c) in the social media domain, and (3) GENIA Kim et al. (2011) in the biomedical domain For unlabeled data, we utilize the training data of each dataset as one data source. For the other data source, we utilize data from external sources, specifically: (1) News Category Dataset (HuffPost) Misra (2022) comprising Huffpost news articles from 2012-2022 for ACE. We filter articles corresponding to political, financial, and business articles, (2) COVIDKB Zong et al. (2022) mining tweets from the Twitter COVID-19 Endpoint released in April 2020 as the external data source, (3) GENIA2013 dataset Kim et al. (2013) as the external data for GENIA. We provide statistics about this data in Table 8.
Data # Sents # Event Average
Source Mentions Length
Test Data
ACE - test 832 403 22.9
SPEED - test 586 672 28.1
GENIA - test 2,151 1,805 29.7
Unlabeled Train Data
ACE - train 17,172 - 15.6
SPEED - train 1,601 - 33.5
GENIA - train 6,431 - 30.1
Unlabeled External Data
HuffPost 43,350 - 17.4
COVIDKB 7,311 - 30.6
GENIA2013 6,542 - 17.4
Table 8: Data Statistics for the various test and unlabeled datasets used in our work.
Appendix C Implementation Details
Here, we provide detailed implementation details for each component and models used in our work. We run most of our experiments on NVIDIA RTX A6000/A100 machines with support for 8 GPUs, while for GPT3.5, we make API calls through OpenAI.
C.1 LLM-based Generation
We provide details about the various hyperparameters for using LLMs in all the components of STAR and FIG. For Llama3-8B-Instruct and Llama3-70B-Instruct, we present the hyperparameters in Table 9; while Table 10 presents the hyperparameters for GPT3.5.
Batch Size 32
Temperature 0.6
Top-p 0.9
Max Generation Length 250
Table 9: Hyperparameters for decoding using Llama3-8B/70B model.
Base LLM gpt-3.5-turbo-0125
Temperature 1.0
Top-p 1.0
Max Generation Length 500
Table 10: Hyperparameters for decoding using GPT3.5 model.
C.2 Few-shot Implementation Details
For the few-shot setting, we can access additional $k$ datapoints per event type to aid better performance. For LLM-based prompting, we simply add these examples in the prompt as in-context examples to help the model do better reasoning/generation. For inverse generation methods (STAR, FIG), we do not add the $k$ triggers to the extracted/generated trigger list, as it leads to a drop in model performance. This can be attributed to the presence of duplicate information as the trigger generation/extraction already accounts for the $k$ triggers. For passage generation/Weak Sup, we append the $k$ datapoints to the synthetically generated data to provide signals from the gold data.
C.3 Downstream Model Training
We choose DEGREE Hsu et al. (2022) as our downstream model for evaluation, a generation-based prompting model that utilizes natural language templates. We implemented the DEGREE model under TextEE framework Huang et al. (2024). Table 11 presents the primary hyperparameters for this model.
Pre-trained LM BART-Large
Training Epochs 25
Warmup Epochs 5
Training Batch Size 32
Eval Batch Size 32
Learning Rate 0.00001
Weight Decay 0.00001
Gradient Clipping 5
Beam Size 1
Negative Samples 15
Max Sequence Length 250
Max Output Length 20
Table 11: Hyperparameters for DEGREE model.
C.4 LLM Fine-tuning
We discuss domain-adapted passage generation through LLM fine-tuning in § 6.4. Specifically, we conduct a low-rank finetuning (LoRA) Hu et al. (2021) to reduce computational overhead to fine-tune Llama3-8B-Instruct. We implement LoRA using the peft and trl packages Mangrulkar et al. (2022); von Werra et al. (2020). We choose the task of causal language modeling (i.e. continual pre-training) to perform domain adaptation on unlabeled in-domain sentences. We utilize cross-entropy loss on the dev split of the unlabeled data to select the best model. We provide additional details about the hyperparameters for this fine-tuning for each dataset in Table 12 below.
ACE
Lora Rank 32
Lora Alpha 16
Lora Dropout 0.1
Learning Rate 0.0001
Weight Decay 0.05
Training Batch Size 32
Training Epochs 3
Eval Steps 20
SPEED
Lora Rank 32
Lora Alpha 16
Lora Dropout 0.1
Learning Rate 0.00008
Weight Decay 0.05
Training Batch Size 32
Training Epochs 10
Eval Steps 20
GENIA
Lora Rank 32
Lora Alpha 16
Lora Dropout 0.1
Learning Rate 0.00008
Weight Decay 0.05
Training Batch Size 32
Training Epochs 6
Eval Steps 20
Table 12: Hyperparameters for LoRA fine-tuning Llama3-8B-Instruct.
Appendix D Additional analyses
In this section, we provide additional analyses to support our main experiments.
Method ACE SPEED GENIA
EI TC EI TC EI TC
STAR 44.9 35.0 21.0 10.1 25.9 19.0
   + mention 44.1 32.9 17.1 10.3 28.7 20.4
   + references 35.5 27.3 19.0 9.2 25.8 18.1
Table 13: Measuring model performance improvement providing domain-specific cues in the form of domain-mention (mention) or domain sentence references (references) to the LLM for STAR. EI: Event Identification F1, TC: Trigger Classification F1.
ACE
A 35-year-old cyclist was hit by a speeding car while riding to work, leaving her with severe injuries, while in a separate incident, a local retail giant filed a petition to restructure its debt, sparking concerns about its financial stability.
As the war on terror raged on, the Mujahideen Advisory Council distributed a statement inviting Arab and foreign media reporters to enter Fallujah and cover the battles, while simultaneously, the ownership of the ancient artifacts was transferred to the museum, with the landlord demanding rent on the premises.
SPEED
As the influencer’s viral challenge went viral, her followers were suddenly struck with a mysterious illness after the splash of a contaminated drink, leading to a shocking explosion of fatalities on social media.
As the community struggled to come to terms with the devastating accident that had claimed the lives of several residents, the authorities swiftly implemented a strict quarantine to prevent the spread of the infectious disease, hoping to mitigate the tragedy.
GENIA
The specific transcription factor was elevated by the presence of the hormone, thereby increasing the expression of the target gene, while the inhibitory protein curbed the activity of a competing transcription factor, preventing the expression of a repressor gene.
The binding of PEBP2/CBF to the promoter region boosts the expression of the gene, which turns on the production of a crucial cytokine in response to the immune response.
Table 14: Example passages of overly long and more stereotypical sentences generated when the domain is mentioned or references are added to the LLM prompt for STAR.
Figure 5: Model performance for FIG as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
D.1 STAR with domain-specific prompt
A simple way to infuse domain specific information in inverse generation pipelines like STAR would be to add domain-related information in the prompts to the LLM. We experiment with two such methods: (1) domain-mention, where we provide the target domain information in the prompt and ask the model to generate accordingly, and (2) domain-reference, where we use some examples from the unlabeled data in the prompt as reference sentences to better guide the passage generation. We provide results for these explorations using the Llama3-8B-Insturct model in Table 13. As observed, the results are generally poor, with an average drop of 0.1-0.6% F1 for domain-mention and 3.1-3.8% F1 for domain-reference. This is majorly because LLMs over-compensate, producing longer and more stereotypical information in their generations which hurts the naturalness of the sentence and causes further domain drift. Furthermore, the LLM makes more errors in mentioning the event as a part of its reasoning, which is utilized to make the generation in the domain style. We provide some qualitative examples for such generations in Table 14. In some ways, it also puts into light and amplifies the gains obtained by doing target domain SFT for FIG as discussed in § 6.4.
Figure 6: Prompt for stage 1 of forward generation for target trigger extraction.
Figure 7: Prompt for stage 2 of forward generation for target trigger extraction.
Figure 8: Prompt for inverse generation for passage generation.
Figure 9: Prompt for forward generation for data refinement.
D.2 Impact of different number of training samples
We conduct a small analysis to study the impact of changing the number of generated samples on the downstream model performance for FIG. We present the results for Llama3-8B-Instruct in Figure 5. As observed, the performance continues to increase as we increase the data from $N=10$ to $N=100$ datapoints per event type. This promises that inverse generation will provide continued improvements by having larger control over data distribution.
Figure 10: Illustration of the interface for the human evaluation of the synthetically generated data. Short instructions are provided at the top. Each query comprises the sentence, annotation, and dataset. The human annotator is expected to score 1-5 for each of the three metrics on the right.
Sentence Score
The sudden crash of the ambulance sent shockwaves through the hospital as medical staff rushed to the scene to monitor the patient’s life signs, but it was too late, as the patient succumbed to the infectious disease. SN: 2 ER: 1 AQ: 1
The wealthy entrepreneur transferred ownership of the struggling tech company to her trusted business partner, relinquishing control and financial responsibility SN: 5 ER: 5 AQ: 5
Taken together, these data suggest that Id1 could be a possible target gene for mediating the effects of BMP-6 in human B cells, whereas Id2 and Id3 not seem to be involved. SN: 4 ER: 3 AQ: 2
Table 15: Illustration examples for the human evaluation metrics. SN: sentence naturalness, ER: event relevance, AQ: annotation quality.
D.3 Human Evaluation Details
We conduct a small human evaluation to judge the quality of the synthetic data in § 6.3. Here, we provide additional details about the human study and evaluation. Since the evaluation is conducted on three diverse and niche domains, we only utilize a single human annotator who is an ED expert and has previously worked on all three datasets as the primary annotator.
We majorly evaluate on three dimensions: (1) Sentence naturalness (SN): this metric judges whether the sentence seems grammatical, natural, and fits the domain of the target data. (2) Event Relevance (ER): this metric is computed only for inverse/hybrid generation methods. This evaluation judges whether the sampled event and trigger are appropriately used to generate a sensible alignment with the target domain. Furthermore, it is verified if the right event definition is used. (3) Annotation Quality (AQ): this metric judges if the right trigger is used for each event mentioned in the synthetic output. If there are any missing events, then this score is penalized. For each metric, a score is given on a Likert scale Likert (1932) from 1 (worst) to 5 (best). We also provide event definitions for each event in each dataset as a reference for better judgment. We illustrate the annotation interface in Figure 10 and provide some sample examples in Table 15.
Dataset Event Method Trigger Sentence
ACE Attack STAR raid As the rebels embarked on a daring trek across the desert, they launched a surprise raid on the heavily guarded fortress, catching the enemy off guard.
siege The government issued fines to the rebel group, while the city was under siege, with gunfire erupting throughout the streets.
slay The wealthy businessman filed a complaint against his former partner, leading to a court proceeding, just as the rival gang leader slays his arch-nemesis in a brutal battle.
FIG shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
bombing As the city was rocked by a devastating bombing, thousands of protesters took to the streets to demand justice and peace.
fight The rival gangs engaged in a brutal fight, sparking a wave of gunfire that left several people injured and buildings damaged.
SPEED Death STAR asphyxiation The hiker’s life was tragically cut short as asphyxiation occurred after she became stuck in the narrow cave crevice.
overdose A young musician’s overdose became the tragic finale of a life cut short in the midst of chaos.
drowning As she struggled to stay afloat, her heart stopped beating, and she succumbed to drowning in the icy waters.
FIG killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
lost As the pandemic spread rapidly across the globe, thousands of people lost their lives due to the deadly virus.
died The elderly man, who had been suffering from a severe case of tuberculosis, died in his sleep.
GENIA Binding STAR merge The regulatory protein’s ability to activate a specific region of the DNA molecule triggers the merge of two proteins, leading to the modification of gene expression.
fuse When the proteins fuse together, the activity of the transcription factor is inhibited, preventing the gene expression from proceeding.
snap When the two proteins snap together, the binding of the complex inhibits the expression of the target gene by deactivating a specific region of the DNA molecule.
FIG bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
translocation The protein translocation to the nucleus triggers the induction of gene expression.
binds When the enzyme binds to the substrate, it activates the addition of a phosphate group to the target molecule, marking a crucial change in its function.
Table 16: Comparison of generated triggers and sentences from STAR and FIG methods
D.4 Additional Qualitative Examples
In § 6.5, we discussed how FIG improves domain drift qualitatively relative to STARand provided some examples. Here, we provide more examples to further support that study in Table 16. This table further demonstrates how STAR can have a domain drift without a lack of domain-specific cues, while FIG is better here.

### v2

@@
-FIG: Forward-Inverse Generation for Low-Resource Domain-specific Event Detection
+SNaRe [Uncaptioned image]: Domain-aware Data Generation for Low-Resource Event Detection
 Abstract
-Event Detection (ED) is the task of identifying typed event mentions of interest from natural language text, which benefits domain-specific reasoning in biomedical, legal, and epidemiological domains.
-However, procuring supervised data for thousands of events for various domains is a laborious and expensive task.
-To this end, existing works have explored synthetic data generation via forward (generating labels for unlabeled sentences) and inverse (generating sentences from generated labels) generations.
-However, forward generation often produces noisy labels, while inverse generation struggles with domain drift and incomplete event annotations.
-To address these challenges, we introduce FIG, a hybrid approach that leverages inverse generation for high-quality data synthesis while anchoring it to domain-specific cues extracted via forward generation on unlabeled target data.
-FIG further enhances its synthetic data by adding missing annotations through forward generation-based refinement.
-Experimentation on three ED datasets from diverse domains reveals that FIG outperforms the best baseline, achieving average gains of 3.3% F1 and 5.4% F1 in the zero-shot and few-shot settings, respectively.
-Analyzing the generated trigger hit rate and human evaluation substantiates FIG’s superior domain alignment and data quality compared to existing baselines.
-We will release our code at https://github.com/PlusLabNLP/FIG.
-FIG: Forward-Inverse Generation for Low-Resource
-Domain-specific Event Detection
+Event Detection (ED) – the task of identifying event mentions from natural language text – is critical for enabling reasoning in highly specialized domains such as biomedicine, law, and epidemiology.
+Data generation has proven to be effective in broadening its utility to wider applications without requiring expensive expert annotations.
+However, when existing generation approaches are applied to specialized domains, they struggle with label noise, where annotations are incorrect, and domain drift, characterized by a distributional mismatch between generated sentences and the target domain.
+To address these issues, we introduce SNaRe, a domain-aware synthetic data generation framework composed of three components: Scout, Narrator, and Refiner.
+Scout extracts triggers from unlabeled target domain data and curates a high-quality domain-specific trigger list using corpus-level statistics to mitigate domain drift.
+Narrator, conditioned on these triggers, generates high-quality domain-aligned sentences, and Refiner identifies additional event mentions, ensuring high annotation quality.
+Experimentation on three diverse domain ED datasets reveals how SNaRe outperforms the best baseline, achieving average F1 gains of 3-7% in the zero-shot/few-shot settings and 4-20% F1 improvement for multilingual generation.
+Analyzing the generated trigger hit rate and human evaluation substantiates SNaRe’s stronger annotation quality and reduced domain drift.
+We will release our code at https://github.com/PlusLabNLP/SNaRe.
+SNaRe [Uncaptioned image]: Domain-aware Data Generation for
+Low-Resource Event Detection
 Tanmay Parekh      Yuxuan Dong      Lucas Bandarkar      Artin Kim I-Hung Hsu      Kai-Wei Chang      Nanyun Peng Computer Science Department, University of California, Los Angeles        Google {tparekh, violetpeng, kwchang}@cs.ucla.edu
 1 Introduction
-Figure 1: Comparing different data generation paradigms.
-Owing to poor LLM ED reasoning skills, forward generation (top) struggles with noisy labels.
-Inverse generation (middle) suffers from domain drift owing to lack of domain-specific cues.
-We introduce FIG (bottom), a hybrid forward and inverse approach, to overcome both issues to generate better quality data.
+Figure 1: Highlighting the errors of existing data generation approaches.
+(a) Using LLMs to generate the labels from in-domain sentences leads to label noise owing to poor LLM reasoning.
+(b) Utilizing LLMs to generate sentences conditioned on event and domain causes domain drift, wherein the synthetic sentence is not aligned with the target domain.
+Finally, in (c), we illustrate how SNaRe minimizes both errors to generate higher quality synthetic data.
 Event Detection (ED) Sundheim (1992); Doddington et al.
 (2004) involves identifying and categorizing significant events from natural language text based on a pre-defined ontology.
-It has applications in widespread domains like biomedicine Pyysalo et al.
+It has widespread applications in domains such as biomedicine Pyysalo et al.
 (2012), epidemiology Parekh et al.
 (2024b, c), law Francesconi et al.
-(2010), etc.
-However, procuring annotations for each domain-specific ontology to train models is expensive and impractical.
-Although recent works introduce zero-shot LLM-reasoning Gao et al.
-(2023); Cai et al.
-(2024), their performance falls short of supervised approaches Huang et al.
 (2010).
-To mitigate the need for extensive annotations to train supervised models, prior work has explored LLM-powered synthetic data generation to enhance downstream model training.
-A widely used approach is a forward generation or weak supervision He et al.
+Due to the high cost of expert-annotated data, synthetic data generation Xu et al.
+(2023) (i.e., generating sentence and event annotations) has emerged as a promising alternative, particularly for practical use-cases in specialized domains.
+However, existing generation approaches often focus on general-domain settings and fail to address the distinct challenges of specialized domains Song et al.
+(2025).
+Weak supervision methods He et al.
 (2021); Chia et al.
-(2022), where labels are assigned to existing unlabeled data (i.e., $X\rightarrow Y^{\prime}$).
-This generation is limited by LLM ED reasoning, often leading to noisy labels.
-To address this, inverse generation has been proposed Meng et al.
-(2022); Wang et al.
-(2023c), where sentences are generated based on sampled or generated labels (i.e $Y\rightarrow X^{\prime}$).
-However, inverse generation introduces a domain drift between synthetic and target data owing to the lack of any domain-specific cues, as illustrated in Figure 1.
-In our work, we propose Forward-Inverse Generation (FIG), a hybrid approach that integrates forward and inverse generation to enhance synthetic data quality.
-Instead of generating event triggers from scratch — an inherently high-variance process — we first apply forward generation to unlabeled data, extracting domain-specific triggers.
-Next, we utilize inverse generation to synthesize diverse sentences conditioned on the domain-specific triggers.
-Finally, a second forward generation annotates any missing events in the generated data to ensure high data quality.
-This approach produces cleaner synthetic data compared to forward generation while maintaining a closer domain alignment with target data than inverse generation, as shown in Figure 1.
-We evaluate FIG on ED datasets from three different domains: ACE Doddington et al.
+(2022) that use LLMs to generate labels for unlabeled sentences frequently introduce label noise (Figure 1(a)), where incorrect or incomplete labels arise due to weak LLM reasoning Huang et al.
+(2024) or limited domain knowledge Song et al.
+(2025).
+Downstream training on such incorrect labels can cause spurious bias propagation.
+Conversely, recent generation approaches Josifoski et al.
+(2023); Ma et al.
+(2024) that utilize LLMs’ self-knowledge to jointly generate labels and sentences struggle with domain drift (Figure 1(b)), often synthesizing sentences that are misaligned with the target domain.
+This can be attributed to the lack of utilization of target domain information for generation and can drastically hamper model training as lexical and structural cues are highly influential for ED Tong et al.
+(2022).
+Overall, label noise and domain drift reduce synthetic data quality, eventually leading to subpar supervised downstream model performance.
+To this end, we propose SNaRe [Uncaptioned image], a novel domain-aware, three-stage data synthesis LLM pipeline comprising the Scout, Narrator, and Refiner modules.
+Scout surveys unlabeled target domain data to identify salient triggers via prompt-based trigger extraction.
+Using corpus-level statistics for automated aggregation and filtering, Scout curates a list of high-quality domain-specific triggers per event type.
+Next, Narrator samples from these domain-specific triggers and utilizes LLMs to synthesize diverse sentences for each event type.
+Utilizing specialized and diverse domain information for conditional generation aids Narrator in generating more domain-aligned sentences, eventually reducing domain drift in the synthesized sentences.
+Since Narrator sentences could mention additional events apart from the input set, we design the Refiner to utilize LLM inference to annotate such missing events in these sentences.
+Narrator’s conditional text generation and Refiner’s missing label annotation aid in reducing the label noise and ensuring high data quality.
+We provide an illustration of SNaRe’s generation in Figure 1(c).
+We benchmark SNaRe on ED datasets from three domains: ACE Doddington et al.
 (2004) (news), SPEED Parekh et al.
-(2024c) (social media), and GENIA2011 Kim et al.
+(2024c) (epidemiology), and GENIA2011 Kim et al.
 (2011) (biomedical).
-Our primary method of evaluation is testing ED performance of DEGREE Hsu et al.
-(2022) trained on the generated data.
-In the zero-shot setting, FIG outperforms STAR Ma et al.
-(2024) (inverse generation baseline) and weak supervision (forward generation baseline) by an average of 16.8% and 3.5% Tri-C F1, respectively.
-Similarly, in the few-shot setting, FIG surpasses the strongest baseline by an average of 6.2% Tri-C F1.
-We study the generated trigger hit rate (relative to the gold trigger set) and demonstrate how FIG’s triggers have about 7.3% better hit rate compared to the best baseline, which contributes to its superior performance.
-Finally, we also conduct human evaluation, which supports FIG’s superior domain relevance and data annotation quality.
+For evaluation, we report the ED performance of DEGREE Hsu et al.
+(2022) trained on the synthesized data.
+Across the zero-shot and few-shot settings, SNaRe performs the best, outperforming the previous state-of-the-art baselines Ding et al.
+(2023); Ma et al.
+(2024) by an average of 3-7% F1 points.
+Under multilingual generation for Arabic and Chinese, SNaRe outshines even more, with improvements of 4-20% F1 over the best baseline.
+Our analysis reveals how SNaRe’s synthesized triggers overlap 4-11% more (relative to baselines) with the gold trigger set, demonstrating the reduction in domain drift.
+Finally, human evaluation provides qualitative evidence for SNaRe’s superior data annotation quality and domain alignment.
+[^1] We will release our code and data upon acceptance.
 2 Problem Definition
@@
 (2004) for this work.
-The task of ED aims to extract mentions of any events of interest from natural language text.
+ED aims to extract mentions of any events of interest from natural language text.
 Following ACE 2005 Doddington et al.
 (2004), we define an event as something that happens or describes a change of state and is labeled by a specific event type.
-The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is referred to as the event mention.
-Event Detection, in particular, requires extracting the event triggers from the sentence and classifying it into one of the pre-defined event types.
-We provide an illustration of this task below where arrested and campaigns are the trigger words for the event types Justice: Arrest-Jail and Conflict: Demonstrate respectively.
+The word/phrase that most distinctly highlights the occurrence of the event is defined as the event trigger, and the trigger-event type pair is known as the event mention.
+Event Detection requires extracting the event triggers from the sentence and classifying them into one of the pre-defined event types.
+We provide an illustration of this task below, where arrested and campaigns trigger the events of Justice: Arrest-Jail and Conflict: Demonstrate, respectively.
 Some 3,000 people have been arrested since the disobedience campaigns began last week.
 […]
-In our work, we specifically focus on ED in diverse and specialized domains (e.g., biomedical), where procuring a training dataset $D_{T}$ of annotated data points is expensive, whereas reasonable-sized unlabeled data $D_{T}^{\prime}$ is available.
+In our work, we specifically focus on ED in diverse and specialized domains (e.g., biomedical), where procuring a training dataset $D_{T}$ of annotated data points is expensive, but unlabeled data $D_{T}^{\prime}$ is available.
 We focus on two realistic low-resource data setups - zero-shot (zero labeled data) and few-shot ($k$ labeled datapoints per event type) settings.
-Different from domain transfer, we do not consider any labeled data for the source domain and directly optimize model performance for the target domain.
-Figure 2: Model Architecture Diagram.
-Top left - illustration of forward generation.
-Bottom left - illustration of inverse generation.
-Right - FIG and its four components.
-We first extract domain-specific triggers via forward generation, then generate passages using inverse generation.
-Forward generation refines missing events, and we sample $N$ data points per event for downstream training.
-3 Methodology
-In this work, we focus on Large Language Model (LLM)-powered synthetic data generation to alleviate the need for domain-specific annotated training data, as zero-shot LLM-based approaches usually struggle in ED in specialized domains Huang et al.
+Unlike domain transfer, we do not consider any labeled data for the source domain, and directly optimize model performance for the target domain.
+3 Related Works
+Figure 2: Model Architecture Diagram highlighting the various components of SNaRe.
+First, Scout extracts and filters domain-specific triggers, then Narrator generates passages conditioned on these triggers.
+Finally, Refiner adds any missing annotations and sample $N$ data points per event for downstream training.
+Data Generation for Information Extraction
+LLM-powered synthetic data generation has been successful for various NLP tasks Li et al.
+(2023b); Wang et al.
+(2023c); Wu et al.
+(2024); Shao et al.
 (2025).
-By generating a large amount of data instances $D_{s}=\{(X,Y)\}$, we can train downstream supervised ED models.
-Here, we first briefly review prior works on forward generation and backward generation.
-Next, we introduce our proposed data generation method FIG.
-3.1 Background
-We focus on two major paradigms for data generation in our work: forward generation and inverse generation. We describe them briefly below and provide illustrations in Figure 2.
-Forward Generation:
-This is a more straightforward manner of data generation, wherein LLMs are utilized to generate labels $Y^{\prime}$ on unlabeled data $X\in D_{T}^{\prime}$ (i.e.
-$X\rightarrow Y^{\prime})$.
-This generation is analogous to weak/distant supervision (Weak Sup) Mintz et al.
+For information extraction, works have explored knowledge retrieval Chen and Feng (2023); Amalvy et al.
+(2023), translation Parekh et al.
+(2024a); Le et al.
+(2024), data re-editing Lee et al.
+(2021); Hu et al.
+(2023), and label extension Zhang et al.
+(2024).
+Recent works utilize LLMs to generate labels for sentences Chia et al.
+(2022); Ye et al.
 (2022); Wang et al.
-(2021); He et al.
-(2021) where noisy/weak labels are assigned to clean sentences and then utilized to train downstream models.
-Consequently, the label quality is dependent on the reasoning capability of the LLM.
-Inverse Generation:
-Inverse generation is a relatively new data generation paradigm Kumar et al.
-(2020); Schick and Schütze (2021).
-This paradigm first generates/samples potential labels $Y$ based on the task definition and then generates plausible $X^{\prime}$ conditioned on the label $Y$ (i.e.
-$Y\rightarrow X^{\prime}$).
-Recent works like SynthIE Josifoski et al.
-(2023) and STAR Ma et al.
-(2024) have explored the usage of LLM-guided inverse generation for information extraction and event extraction tasks.
-Inverse generation provides control over data distribution via prompting, but the data quality is dependent on the LLM’s sentence generation capability.
-3.2 FIG
-LLM reasoning has been poor for ED Li et al.
-(2023a); Huang et al.
-(2024), which makes forward generation vulnerable to noisy label quality.
-Since LLMs are pre-trained to generate natural sentences, their generation capabilities are stronger, which supports inverse generation.
-However, owing to the lack of any domain-specific information, inverse generation can synthesize highly diverse sentences, leading to a domain drift.
-Merely specifying the domain in the prompt is not sufficient (shown in § D.1).
-Furthermore, inverse generation sentences can mention additional events which remain unannotated introducing noise in the data.
-To this end, we propose Forward-Inverse Generation (FIG), a hybrid approach that infuses forward generation reasoning into the base pipeline of inverse generation to ensure closer alignment to the target domain while improving the quality of the annotated labels.
-To procure domain-specific cues, we assume access to an unlabeled target domain data $D_{T}^{\prime}$ (similar to forward generation).
-We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
-Target Structure Extraction:
-Similar to inverse generation, we first create potential triggers (i.e.
-labels) for each event type.
-However, generation from scratch can lead to very divergent set of triggers.
-For example, an “Attack" event can refer to a war in the news, cyberattacks in the cybersecurity domain, diseases in epidemiology, or criticism in economics.
-Naturally, each scenario would assume different triggers, which inverse generation struggles to generate even when provided with clear event definitions (actual examples in § 6.5).
-Instead, we generate potential triggers per event type by utilizing an LLM to extract triggers using a forward generation over the unlabeled data $D_{T}^{\prime}$.
-To ensure highly precise extraction of triggers, we develop a two-stage prompt setup.
-The first stage is tasked to identify and filter possible event types mentioned in the text based on the task instructions and event definitions.
-The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type.
-We illustrate this setup in § A.
-After extraction, we aggregate and sort the triggers for each event type at the corpus level and filter out the top $t$ triggers for each event type as the set of clean triggers.
-This filtering helps remove many noisy triggers introduced in the forward generation.
-Finally, we create target structures $Y$ by sampling 1-2 event types per instance and sampling triggers from the clean trigger set for each sampled event type.
-We ensure uniform sampling of event types and triggers to avoid unbalanced label distribution.
-Overall, such forward generation-based trigger extraction ensures the distillation of domain-specific knowledge in the target structures.
-Passage Generation:
-This component is the core of inverse generation and is tasked with generating passages corresponding to the constraints of the sampled target structure.
-Specifically, the LLM is prompted with the task instructions, the event definitions, the sampled target structure $Y$, and asked to generate a passage $X^{\prime}$ which mentions the event types using the triggers in $Y$.
-We illustrate this prompt in Figure 8.
-The key to generating better target data aligned passages lies in the presence of the target data aligned target structure $Y$ (shown via examples in § 6.5).
-Data Refinement and Sampling:
-While passage generation ensures the target structure events are mentioned in the passage, the passage could still have other unknown events which can lead to under-annotated $(X^{\prime},Y)$ data instances.
-We provide an illustration in Figure 3 with the target trigger positive, but the generated sentence also has a missing mention of symptom event triggered by got.
-To account for such missing annotations, we utilize forward generation to identify all possible event mentions in the generated passage.
-We remove duplicates and add the new event mentions to create a more complete target structure $Y_{f}$.
-Missed Event Trigger: “got” Event: symptom Generation Trigger: “Positive” Event: infect Ok, I just got a fever… Theres a possibility Im COVID-19 Positive
-Figure 3: Illustration of how inverse generation can produce unannotated event mentions. Blue box = target event mention, red box = unannotated event mention.
-To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger.
-Additionally, we standardize trigger annotations by correcting variations in trigger word forms.
-Finally, we apply a greedy sampling algorithm to sample $N$ instances $(X^{\prime},Y_{f})$ per event type to create our final synthetic dataset $D_{s}$.
-Downstream Model Training:
-The final component utilizes the generated synthetic data $D_{s}$ to train downstream ED models in a supervised manner.
-The trained ED models are then used for inferring on the test set and for eventual evaluation.
-4 Experimental Setup
-Base LLM Method Unlabeled ACE SPEED GENIA Average
-Data Source Eve-I Tri-C Eve-I Tri-C Eve-I Tri-C Eve-I Tri-C
-Llama3-8B Inference - 30.2 23.8 39.8 25.4 21.9 17.2 30.6 22.1
-STAR - 44.9 35.0 21.0 10.1 25.9 19.0 30.6 21.4
-Weak Sup train 41.7 37.8 45.6 31.5 26.9 21.4 38.1 30.2
-FIG(ours) train 57.4 50.2 44.6 31.5 35.2 28.9 45.7 36.9
-FIG (ours) external 57.7 52.6 47.8 32.9 33.6 24.6 46.4 36.7
-Llama3-70B Inference - 46.9 41.3 46.9 35.6 34.2 28.2 42.7 35.0
-STAR - 50.0 42.3 18.3 13.8 23.3 16.9 30.5 24.3
-Weak Sup train 53.2 48.0 52.8 39.6 36.2 29.1 47.4 38.9
-FIG (ours) train 58.1 53.8 49.9 38.7 38.0 29.7 48.7 40.7
-FIG (ours) external 59.7 55.6 50.1 39.2 39.2 31.5 49.7 42.1
-GPT-3.5 Inference - 33.0 26.2 44.2 32.9 31.2 24.7 36.1 27.9
-STAR - 45.0 36.6 21.3 14.6 21.8 14.3 29.4 21.8
-Weak Sup train 49.7 44.6 50.7 37.5 37.7 30.1 46.1 37.4
-FIG (ours) train 54.8 48.3 50.3 36.8 39.3 31.1 48.1 38.7
-FIG (ours) external 54.0 48.5 50.1 36.1 38.7 29.4 47.6 38.0
-- Gold Data - 64.6 61.6 64.0 53.5 51.3 44.0 60.0 53.0
-Table 1: Zero-shot results comparing FIG with other baselines across three datasets and three base LLMs.
-Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al.
-(2022) model trained on data generated by each technique.
-Eve-I: Event Identification F1, Tri-C: Trigger Classification F1.
-Datasets:
-We consider three ED datasets from diverse domains for our experiments: (1) ACE Doddington et al.
-(2004), in the news domain, (2) SPEED Parekh et al.
-(2024c), in the social media domain, and (3) GENIA Kim et al.
-(2011), in the biomedical domain.
-We simplify GENIA by converting the original document level annotations to sentence-level annotations.
-For the few-shot setting, we compile $k$ data instances from the training data as the few-shot examplers.
-For our unlabeled data, we consider two sources: (1) Train - annotation-free training splits (i.e.
-only the text) of each dataset and (2) External - unlabeled data from other external sources.
-For ACE, we utilize News Category Dataset Misra (2022) comprising Huffpost news articles from 2012-2022 as the external data source.
-We filter articles corresponding to political, financial, and business articles.
-For SPEED, we utilize COVIDKB Zong et al.
-(2022) mining tweets from the Twitter COVID-19 Endpoint released in April 2020 as the external data source.
-Finally, we utilize GENIA2013 dataset Kim et al.
-(2013) as the external data for GENIA.
-We provide statistics about these datasets in Table 8.
-Baseline methods:
-We consider three LLM-based techniques for low-resource ED as the baselines for our work.
-(1) Inference Gao et al.
-(2023): LLMs are used to directly infer on the target test data using their reasoning capability.
-(2) STAR Ma et al.
-(2024): This model is the state-of-the-art inverse generation model for ED.
-It utilizes trigger generation, passage generation, and data refinement steps without using any unlabeled data, (3) Weak Supervision (Weak Sup): This model acts as the forward generation baseline.
-We utilize the Inference model to provide labels for the unlabeled data.
-For an upper bound reference, we also include a human generation baseline (Human) wherein we sample $N$ data instances from the gold training data of each dataset to train the downstream ED model.
-Base models:
-For our base LLMs, we consider three instruction-tuned LLMs of varying sizes, namely Llama3-8B-Instruct (8B model), Llama3-70B-Instruct (70B model) Dubey et al.
-(2024), and GPT-3.5 (175B model) Brown et al.
-(2020).
-For our downstream ED model, we consider a specialized low-resource model DEGREE Hsu et al.
-(2022), a generative model prompted to fill event templates powered on a BART-large pre-trained language model Lewis et al.
-(2020).
-Evaluation:
-Our primary evaluation metric is the synthesized data-trained model’s ED performance on the final test splits of each dataset.
-We consider two low-resource settings - zero-shot (no labeled data) and few-shot ($k$ datapoints per event type are used).
-Note this is different cross-dataset works Cai et al.
-(2024) which train and test on an exclusive set of event types.
-For Inference, the LLM is directly run on the test set to procure model predictions.
-We report the F1 scores for two metrics Ahn (2006) for measuring model performance: (1) Event Identification (Eve-I) - correct identification of events, and (2) Trigger Classification (Tri-C) - correct identification of trigger-event pairs.
-Implementation Details:
-We follow STAR for the implementation of the baseline models and the majority of hyperparameter settings.
-For FIG’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation.
-We generate $N=50$ datapoints per event type for each generation strategy.
-All our experimental results are reported over an average of three runs.
-Additional implementation details are provided in Appendix C.
-Figure 4: Few-shot results comparing FIG with other baselines across three datasets using Llama3-8B-Instruct as the base LLM.
-Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al.
-(2022) model trained on data generated by each technique.
-Tri-C: Trigger Classification F1, #: Number of.
-5 Results
-We present the results and insights for our zero-shot and few-shot settings below.
-5.1 Zero-shot Results
-We present the main results comparing the various methods across different LLMs and datasets in Table 1.
-For FIG, we show results using both the train data as well as the external data as the external data source.
-We highlight some of our main findings below.
-Forward generation > Inverse generation:
-Although STAR performs decently for ACE, its performances for SPEED and GENIA are quite poor (even below the Inference baseline).
-Since SPEED (social media) and GENIA (biomedical) are more domain-specific than ACE (news), we attribute this poor performance to STAR’s inability to generate domain-specific data instances (further validated in § 6.3).
-On the other hand, Weak Sup outperforms STAR, providing average gains of 13% Tri-C F1, thus establishing the superiority of forward generation-based synthetic data.
-FIG performs the best:
-Our hybrid approach combines the benefits from both forward and inverse generation and provides the best performance.
-On average, FIG beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points - proving how domain-specific cues from unlabeled data can help generate better-aligned inverse generated data.
-Compared to forward generation Weak Sup baseline, FIG provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1 - suggesting how data diversity and cleaner label quality can help improve model performance.
-External data source is effective:
-Assuming access to the training data as the unlabeled data source can be a strong assumption and bias for FIG.
-To cross-validate this assumption’s impact, we also consider FIG with external data sources.
-Surprisingly, as seen from Table 1, FIG with external data provides similar gains of 4% Eve-I F1 and 3.4% Tri-C F1 over the best baseline.
-This demonstrates that instead of getting biased by external data, FIG utilizes the useful domain-specific signals (in the form of extracted triggers) from the external data and provides similar/better gains as with using the training data.
-Generation ability doesn’t scale up as reasoning ability:
-Although FIG performs better across all LLMs, the gains are much higher for Llama3-8B model relative to the Llama3-70B or GPT3.5 models.
-Comparing the improvements in model performance when scaling up from Llama3-8B to LLama3-70B, inverse generation-centric methods (STAR, FIG) improve by an average of 2.5% F1 points while reasoning-centric methods (Inference, Weak Sup) improve considerably better by an average of 10.8% F1 points.
-A similar disparity in performance improvements is observed for GPT3.5 as well.
-This sheds light on how the generation capabilities of LLMs do not scale compared to reasoning capabilities.
-Simultaneously, it shows that FIG is particularly effective when used with smaller LLMs whose the reasoning capability are poor.
-5.2 Few-shot Results
-We also study the various methods in the presence of small annotated data as part of our few-shot experiments.
-Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized.
-Majorly, we utilize the $k$ shots as in-context examples in the LLM prompts where applicable and append these few-shot examples to the synthesized training data as well.
-Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$ shot examples.
-We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 4 and summarize our major findings below.
-FIG consistently performs the best:
-Similar to zero-shot results, we observe that FIG consistently beats all other baseline models. On an average, FIG outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
-Generation abilities improve significantly:
-Contrary to our findings of small improvements when scaling up LLM model size in § 5, we observe that inverse generation-centric methods (STAR, FIG) consistently improve as we increase the number of shots.
-Contrastingly, reasoning-centric models stagnate and don’t improve as much.
-This highlights how inverse generation becomes particularly effective in the presence of few exemplars.
-6 Analysis
-In this section, we provide various analyses to study FIG’s superior performance. Unless specified, we utilize Llama3-8B-Instruct as the base LLM for the analyses.
-Method ACE SPEED GENIA
-FIG 50.2 31.5 28.9
-   – Trigger Generation 43.2 27.8 28.2
-   – Data Refinement 47.4 23.3 22.0
-Table 2: Ablation study for FIG’s various components measured as Tri-C F1 performance across datasets.
-Method ACE SPEED GENIA
-EI TC EI TC EI TC
-FIG 57.4 50.2 44.6 31.5 35.2 28.9
-Weak Sup + STAR 46.9 38.9 44.5 29.5 30.2 24.3
-Table 3: Comparing FIG’s hybrid approach with data-mixing of forward and inverse generation.
-Method ACE SPEED GENIA
-STAR 9.6% 15.3% 15.1%
-Weak Sup 19.1% 38.4% 44.8%
-FIG 23.2% 49.4% 52.5%
-Table 4: Reporting the hit rate of synthetic data triggers relative to gold test triggers for the three methods.
-6.1 Ablation study
-Table 2 shows the ablation study for each of our forward generation components.
-We observe how forward generation is critical in both trigger generation and data refinement stages with average performance reductions of 3.8% F1 and 6% F1 on removing the respective components from FIG.
-6.2 Comparison with Data mixing
-Data-mixing Hoffmann et al.
-(2022); Xie et al.
-(2023) is a widely used technique for leveraging complementary information across datasets to promote robust downstream model training.
-We mix forward (Weak Sup) and inverse generation (STAR) as a hybrid baseline to compare with FIG.
-To keep comparisons fair, we consider $N/2$ data instances per event type from each dataset.
-Results from Table 3 demonstrate how FIG beats the data-mixing based hybrid model by 5-6% F1, highlighting the need for explicit model design to combine the benefits of forward and inverse generation.
-6.3 Analyzing trigger quality and drift
-ED models have a strong tendency to over-rely on lexical relations between triggers and events Tong et al.
+(2023a); Tang et al.
+(2023), while some other works explore the generation of sentences from labels Josifoski et al.
 (2023); Ma et al.
 (2024).
-Thus, we compare the synthetic data triggers with the gold test triggers as a raw study on the quality and drift of triggers in the synthesized data.
-Specifically, we extract triggers per event type in both datasets and measure the hit rate of the synthesized triggers on the gold set, as reported in Table 4.
-As observed, the poor hit rate shows the poor overlap of STAR’s triggers with the gold triggers, which is a primary reason for its domain drift.
-Furthermore, the consistently stronger precision of FIG relative to Weak Sup demonstrate FIG’s better trigger annotation quality.
-Method Naturalness Event Annotation
-Relevance Quality
-STAR 3.1 3.4 3.1
-Weak Sup 4.2 - 2.9
-FIG 3.6 4.0 3.6
-Table 5: Human evaluation for sentence naturalness, relevance of event in generated sentence, and the annotation quality. 1 = worst, 5 = best.
-Dataset Event Method Trigger Sentence
-ACE Attack STAR raid As the rebels embarked on a daring trek across the desert, they launched a surprise raid on the heavily guarded fortress, catching the enemy off guard.
-FIG shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
-SPEED Death STAR asphyxiation The hiker’s life was tragically cut short as asphyxiation occurred after she became stuck in the narrow cave crevice.
-FIG killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
-GENIA Binding STAR merge The regulatory protein’s ability to activate a specific region of the DNA triggers the merge of two proteins, leading to the modification of gene expression.
-FIG bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
-Table 6: Qualitative examples demonstrating STAR and FIG’s trigger and sentence generation quality.
-Method ACE SPEED GENIA
-EI TC EI TC EI TC
-FIG 57.4 50.2 44.6 31.5 35.2 28.9
-   + SFT LLM 55.2 51.7 46.9 35.8 36.7 29.1
-Table 7: Measuring model performance improvement using a LLM fine-tuned on unlabeled train data (SFT LLM) for FIG. EI: Event Identification F1, TC: Trigger Classification F1.
-To further study the label quality and relevance, we conduct a human evaluation.
-Specifically, a human expert ED annotator is tasked to score the generations (between 1-5) on the naturalness of the sentence for the domain, the correct relevance of the event mentioned in the generated sentence, and the annotation quality (more details in § D.3).
-We provide the averaged scores across the three datasets for 90 samples in Table 5.
-Weak Sup has high sentence quality but poor label annotations; while STAR suffers from poor event relevance indicating domain drift.
-Overall, FIG performs the best with high annotation quality and event relevance.
-6.4 Domain-adapted LLM Fine-tuning
-We fine-tune the base LLM used for passage generation on the unlabeled target-domain data $D_{T}^{\prime}$ to better align the generated passages to the target domain.
-Naturally, this can be applied only for smaller LLMs owing to fine-tuning costs.
-We present the results of fine-tuning Llama3-8B-Instruct on the unlabeled train data in Table 7.
-On average, we observe how target data fine-tuning improves FIG by a slight 0.5-2% F1, suggesting that target-domain passage generation may help, but it is not a highly influencing factor to improve downstream model performance.
-6.5 Qualitative analysis of generated data
-We provide qualitative evidence for FIG’s domain-aligned triggers compared to STAR in Table 6 (more examples in Table 16).
-Owing to lack of domain grounding in STAR, the resulting triggers often appear misaligned (e.g.
-asphyxiation for death event related to pandemics).
-This misalignment carries over to the generated sentences, further reducing their quality and alignment.
-In contrast, FIG’s triggers are better aligned to the target domain corpus, resulting in better quality data.
-7 Related Works
+Our work introduces SNaRe focused on infusing better domain-specific information for data generation.
 Low-resource Event Detection
@@
 (2022).
-To address the growing need for event detection across expanding domains, better low-resource domain-specific techniques are essential.
-Prior works have explored transfer learning via Abstract Meaning Representation Huang et al.
+To address the growing need for event detection across expanding domains, prior works have explored transfer learning via Abstract Meaning Representation Huang et al.
 (2018), Semantic Role Labeling Zhang et al.
@@
 Reformulating ED as a conditional generation task has also aided low-resource training Hsu et al.
-(2022, 2023); Huang et al.
+(2022, 2023b); Huang et al.
 (2022).
@@
 (2023b) and transfer-learning Cai et al.
-(2024) has been explored for low-resource ED and transfer learning, but performance remains inferior to supervised models Huang et al.
+(2024) has been explored, but their performance remains inferior to supervised models Huang et al.
 (2024).
-This motivates efforts in synthetic data generation for low-resource ED.
-Data Generation for Information Extraction
-LLM-powered synthetic data generation has been successful for various NLP tasks Li et al.
-(2023b); Wang et al.
-(2023c); Wu et al.
-(2024); Shao et al.
+This motivates efforts in LLM-powered synthetic data generation for low-resource ED.
+4 Methodology - SNaRe
+In this work, we focus on domain-aware synthetic data generation using Large Language Models (LLMs) to alleviate the need for expert-annotated training data.
+By generating a large dataset $D_{s}=\{(X,Y)\}$, we can train domain-specific downstream ED models using minimal supervised data.
+Existing weak-supervision approaches Mintz et al.
+(2009); Wang et al.
+(2021) utilizing automatic methods to assign labels to unlabeled sentences often generate incorrect labels (label noise).
+This can be attributed to the domain-specific context understanding and deep reasoning requirement of ED, leading to poor automatic label quality, even when using recent LLMs Huang et al.
 (2024).
-For information extraction, works have explored knowledge retrieval Chen and Feng (2023); Amalvy et al.
-(2023), translation Parekh et al.
-(2024a); Le et al.
-(2024), and data re-editing Lee et al.
-(2021); Hu et al.
+Another route of approaches that utilize LLMs to generate sentences conditioned on labels Schick and Schütze (2021); Josifoski et al.
+(2023), i.e., synthesize $X$ for a designated $Y$, often curate sentences that are distributionally divergent from the target domain (domain drift).
+This is mainly since these approaches focus on the general domain and fail to utilize any target domain signals in their generation.
+Both label noise and domain drift hurt the synthetic data quality, in turn, diminishing the downstream supervised model performance.
+To mitigate these issues, we propose SNaRe [Uncaptioned image], a domain-aware high-quality data synthesizer, comprising three components: Scout, Narrator, and Refiner.
+Scout studies unlabeled target domain data $D_{T}^{\prime}$ to curate domain-specific triggers, in turn, reducing domain drift.
+Narrator generates domain-specific sentences conditioned on Scout’s curated triggers, while Refiner adds additional annotations to ensure high-quality labels.
+Overall, SNaRe is a training-free LLM inference pipeline and easily deployable and scalable across domains.
+We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
+4.1 Scout
+Scout is tasked with the curation of domain-specific triggers that are later utilized for sentence generation.
+Events can assume a wide range of triggers depending on the domain and context.
+For example, an "Attack" event can be triggered by war, killed in news, or breach, phish in cybersecurity, or infect, transmit in epidemiology domains.
+Thus, unlike past works Ma et al.
+(2024) that utilize only LLMs’ internal knowledge for trigger generation, we develop Scout, which extracts high-precision domain-specific triggers using unlabeled target domain data $D_{T}^{\prime}$.
+Specifically, trigger extraction involves a two-stage prompt setup: (1) The first stage is tasked with identifying and filtering possible event types mentioned in the target domain sentence, and (2) The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type.
+We illustrate this setup in § A.
+To ensure high-precision of the triggers, we develop an aggregation and filtering mechanism by incorporating corpus-level statistics.
+Specifically, for each event type, we aggregate the counts of the triggers at the corpus level and filter out the top $t=10$ triggers as the curated list of high-quality domain-specific event-indexed triggers.
+These triggers carry important target domain signals that help in generating domain-specific sentences (§ 4.2), in turn reducing domain drift of our synthetic data.
+4.2 Narrator
+Narrator is tasked with the synthesis of domain-specific sentences for our synthetic dataset.
+Existing works Josifoski et al.
+(2023); Ma et al.
+(2024) do not utilize any target domain information, which causes domain drift in their synthesized sentences.
+Instead, in our work, we condition the Narrator to utilize the rich and diverse domain-specific triggers from Scout to synthesize domain-specific sentences, which, in turn, reduce domain drift.
+Specifically, Narrator samples 1-2 event types per synthetic data instance and corresponding domain-specific triggers from Scout’s curated trigger list – constituting the label $Y$.
+Next, it prompts the LLM with the task instructions and the event definitions, and asks it to generate a passage $X^{\prime}$ that mentions the sampled events using the sampled triggers ($Y$).
+We illustrate this prompt in Figure 8.
+The generated sentences are naturally more aligned with the target domain owing to the conditioning on the domain-specific triggers (qualitative examples shown in § 7.6).
+Further domain-specific adaptations are possible by fine-tuning the LLM on unlabeled domain-specific data (analysis in § 7.5).
+Base LLM Method Unlabeled ACE SPEED GENIA Average
+Data Source Eve-I Tri-C Eve-I Tri-C Eve-I Tri-C Eve-I Tri-C
+Llama3-8B Inference - 30.2 23.8 39.8 25.4 21.9 17.2 30.6 22.1
+STAR - 44.9 35.0 21.0 10.1 25.9 19.0 30.6 21.4
+Weak Sup train 41.7 37.8 45.6 31.5 26.9 21.4 38.1 30.2
+SNaRe (ours) train 57.4 50.2 44.6 31.5 35.2 28.9 45.7 36.9
+SNaRe (ours) external 57.7 52.6 47.8 32.9 33.6 24.6 46.4 36.7
+Llama3-70B Inference - 46.9 41.3 46.9 35.6 34.2 28.2 42.7 35.0
+STAR - 50.0 42.3 18.3 13.8 23.3 16.9 30.5 24.3
+Weak Sup train 53.2 48.0 52.8 39.6 36.2 29.1 47.4 38.9
+SNaRe (ours) train 58.1 53.8 49.9 38.7 38.0 29.7 48.7 40.7
+SNaRe (ours) external 59.7 55.6 50.1 39.2 39.2 31.5 49.7 42.1
+GPT-3.5 Inference - 33.0 26.2 44.2 32.9 31.2 24.7 36.1 27.9
+STAR - 45.0 36.6 21.3 14.6 21.8 14.3 29.4 21.8
+Weak Sup train 49.7 44.6 50.7 37.5 37.7 30.1 46.1 37.4
+SNaRe (ours) train 54.8 48.3 50.3 36.8 39.3 31.1 48.1 38.7
+SNaRe (ours) external 54.0 48.5 50.1 36.1 38.7 29.4 47.6 38.0
+(Upper Bound) Gold Data - 64.6 61.6 64.0 53.5 51.3 44.0 60.0 53.0
+Table 1: Zero-shot results comparing SNaRe with other baselines across three datasets and three base LLMs.
+Except for Inference, all other evaluations are performances of downstream DEGREE Hsu et al.
+(2022) model trained on data generated by each technique (50 datapoints per event type).
+4.3 Refiner
+While the Narrator ensures the sampled triggers are mentioned in the passage, it could potentially introduce new unknown events in the passage, leading to under-annotated $(X^{\prime},Y)$ data instances.
+We illustrate this in Figure 3 where the sampled trigger positive for infect event is mentioned, but the sentence also mentions the missing symptom event triggered by got.
+To account for such missing annotations, we introduce Refiner - tasked to annotate missing events in the generated sentence.
+Here, we simply prompt the LLM to find all event mentions in the generated sentence (illustrated in Figure 9) and append them to the original sampled $Y$.
+To avoid adding noisy labels, we only add annotations for newly discovered events.
+To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger.
+Additionally, we standardize trigger annotations by correcting variations in trigger word forms.
+Such normalization and conservative filtering further aid in ensuring high data quality.
+Finally, we apply a greedy sampling algorithm to sample $N=50$ instances $(X^{\prime},Y)$ per event type to create our final synthetic dataset $D_{s}$.
+Missed Event Trigger: “got” Event: symptom Generation Trigger: “Positive” Event: infect Ok, I just got a fever… Theres a possibility Im COVID-19 Positive
+Figure 3: Illustration of how inverse generation can produce unannotated event mentions. Blue box = target event mention, red box = unannotated event mention.
+Downstream Model Training:
+The final component utilizes the generated synthetic data $D_{s}$ to train downstream ED models in a supervised manner.
+The trained ED models are then used to infer on the test set and for eventual evaluation.
+Since we use small BART-based language models for inference, our inference time computation is negligible compared to LLM inference methods.
+5 Experimental Setup
+Datasets:
+We consider three ED datasets from diverse domains for our experiments: (1) ACE Doddington et al.
+(2004), in the news domain, (2) SPEED Parekh et al.
+(2024c), in the social media domain, and (3) GENIA Kim et al.
+(2011), in the biomedical domain.
+We simplify GENIA by converting the original document-level annotations to sentence-level annotations.
+We consider the Arabic and Chinese versions of ACE for cross-lingual experiments.
+For the few-shot setting, we sample $k$ few-shot examples from the training data.
+For our unlabeled data, we consider two sources: (1) Train - annotation-free training splits (i.e., only the text) of each dataset and (2) External - unlabeled data from other external sources.
+For this external data source, for ACE, we utilize News Category Dataset Misra (2022) comprising Huffpost news articles from 2012-2022.
+We filter articles corresponding to political, financial, and business articles.
+For SPEED, we utilize COVIDKB Zong et al.
+(2022), comprising tweets from the Twitter COVID-19 Endpoint.
+Finally, we utilize GENIA2013 Kim et al.
 (2013).
-Some works in the directions we focus in our work include forward generation Chia et al.
-(2022); Ye et al.
-(2022); Wang et al.
-(2023a); Tang et al.
-(2023) and inverse generation Josifoski et al.
-(2023); Ma et al.
+We provide statistics about these datasets in Table 10.
+Baseline methods:
+We consider three LLM-based techniques for low-resource ED as the baselines for our work.
+(1) Inference Gao et al.
+(2023): LLMs are used to directly infer on the target test data using their reasoning capability.
+(2) STAR Ma et al.
+(2024): The state-of-the-art generation model for ED utilizing LLMs for trigger and passage generation without using any unlabeled data, (3) Weak Supervision (Weak Sup) Ding et al.
+(2023): LLMs are utilized to synthesize labels for unlabeled data.
+For an upper bound reference, we also include a Gold Data generation baseline wherein we sample from the gold training data of each dataset to train the downstream ED model.
+Base models:
+For our base LLMs, we consider three instruction-tuned LLMs of varying sizes, namely Llama3-8B-Instruct (8B model), Llama3-70B-Instruct (70B model) Dubey et al.
+(2024), and GPT-3.5 (175B model) Brown et al.
 (2020).
-We improve on existing forward and inverse generation works via our hybrid generation approach FIG.
+For our downstream ED model, we consider a specialized low-resource model DEGREE Hsu et al.
+(2022), a generative model prompted to fill event templates powered by a BART-large pre-trained language model Lewis et al.
+(2020).
+Figure 4: Few-shot results comparing SNaRe with other baselines across three datasets using Llama3-8B-Instruct as the base LLM.
+Except for Inference, all other evaluations are performances of the downstream DEGREE Hsu et al.
+(2022) model trained on data generated by each technique.
+Tri-C: Trigger Classification F1, #: Number of.
+Evaluation:
+Our primary evaluation metric is supervised model performance trained on the synthesized data.
+We consider two low-resource settings - zero-shot (no labeled data) and few-shot ($k$ datapoints per event type are used).
+For Inference baseline, the LLM is directly run on the test set to procure model predictions.
+We report the F1 scores for two metrics Ahn (2006): (1) Event Identification (Eve-I) - correct identification of events, and (2) Trigger Classification (Tri-C) - correct identification of trigger-event pairs.
+Implementation Details:
+We follow STAR for the implementation of the baseline models and most hyperparameter settings.
+For SNaRe’s passage generation, we select the top $t=10$ triggers (except $t=8$ for GENIA) for passage generation.
+We generate $N=50$ datapoints per event type for each generation strategy.
+All our experimental results are reported over an average of three runs.
+Additional details are provided in Appendix C.
+6 Results
+We present the results for our zero-shot, few-shot settings, and cross-lingual experiment below.
+6.1 Zero-shot Results
+We present the main zero-shot results comparing the baselines across different LLMs and datasets in Table 1 and discuss our findings below.
+SNaRe performs the best:
+On average, SNaRe beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points – demonstrating how domain-specific cues from unlabeled data aid in tackling domain drift.
+Compared to Weak Sup, SNaRe provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1, suggesting how cleaner label quality can help improve model performance.
+External data source is effective:
+Assuming access to the training data as the unlabeled data source can be a strong assumption and bias for SNaRe.
+To verify the robustness of our approach, we also evaluate SNaRe with external data sources.
+Surprisingly, as seen in Table 1, SNaRe with external data provides similar gains of 4% Eve-I F1 and 3.4% Tri-C F1 over the best baseline.
+We posit that the higher volume of external data leads to the extraction of cleaner domain-specific triggers, which eventually aids in better downstream performance.
+LLM Method Arabic Chinese
+EI TC EI TC
+Llama3-8B Inference 21.5 13.4 15.0 11.8
+STAR 11.5 10.5 19.7 16.0
+Weak Sup 21.5 16.2 26.3 19.1
+SNaRe (ours) 40.1 33.6 35.9 31.1
+Llama3-70B Inference 37.5 27.7 32.0 29.1
+STAR 37.1 30.9 26.0 22.0
+Weak Sup 30.0 20.4 28.1 26.3
+SNaRe (ours) 47.5 44.0 40.4 33.1
+Table 2: Comparing SNaRe with zero-shot inference for other powerful LLMs for the ACE dataset.
+6.2 Few-shot Results
+We also study the various methods in the presence of small annotated data as part of our few-shot experiments.
+Specifically, we study the $k=2$ and $k=5$ few-shot settings, where $k$ annotated examples per event type are utilized.
+We utilize the $k$-shots as in-context examples in the LLM prompts and append these few-shot examples to the synthesized training data as well.
+Additionally, we consider another baseline (Supervised) of downstream models trained only on the $k$-shot examples.
+We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 4.
+Similar to zero-shot results, we observe that SNaRe consistently beats all other baseline models.
+On average, SNaRe outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
+6.3 Zero-shot Multilingual Results
+To highlight the utility of our work, we apply our work across languages, specifically Arabic (ar) and Chinese (zh).
+We used multilingual ACE data Doddington et al.
+(2004) for this experiment and utilized TagPrime Hsu et al.
+(2023a), powered by XLM-Roberta-large Conneau et al.
+(2019), as the downstream ED model.
+We present our results for Llama3-8B-Instruct and Llama3-70B-Instruct LLMs in Table 2.
+Surprisingly, SNaRe performs the best out-of-the-box, with improvements ranging 10-20% F1 for Arabic and 4-12% F1 for Chinese, highlighting the broader impact of our work.
+7 Analysis
+In this section, we study the superior performance of SNaRe through various analyses. Unless specified, we use Llama3-8B-Instruct as the base LLM.
+Method ACE SPEED GENIA
+SNaRe 50.2 31.5 28.9
+   – Scout 43.2 27.8 28.2
+   – Refiner 47.4 23.3 22.0
+Table 3: Ablation study for SNaRe’s Scout and Refiner measured as Tri-C F1 performance across datasets.
+LLM + Method Eve-I Tri-C
+Llama3-70B + Inference 46.9 41.3
+Llama3.3-70B + Inference 48.9 43.0
+Qwen2.5-72B + Inference 40.9 34.2
+GPT4o-mini + Inference 34.5 28.8
+GPT4o + Inference 51.4 47.7
+QwQ-32B + Inference 49.7 43.5
+Deepseek-R1-L3-70B + Inference 41.8 36.6
+Llama3-70B + SNaRe (train) 58.1 53.8
+Llama3-70B + SNaRe (external) 59.7 55.6
+Table 4: Comparing SNaRe with zero-shot inference for other powerful LLMs for the ACE dataset.
+Method ACE SPEED GENIA
+EI TC EI TC EI TC
+SNaRe 57.4 50.2 44.6 31.5 35.2 28.9
+Weak Sup + STAR 46.9 38.9 44.5 29.5 30.2 24.3
+Table 5: Comparing SNaRe with data-mixing of synthesized data from previous works.
+Method ACE SPEED GENIA
+STAR 9.6% 15.3% 15.1%
+Weak Sup 19.1% 38.4% 44.8%
+SNaRe 23.2% 49.4% 52.5%
+Table 6: Reporting the hit rate of synthesized data triggers relative to gold test triggers.
+7.1 Ablation study
+Table 3 shows the ablation study for Scout and Refiner.
+For ablating Scout, we replace it by prompting LLM to directly generate triggers.
+We observe how both Scout and Refiner are critical, with average performance reductions of 3.8% F1 and 6% F1 upon removing the respective components.
+Dataset Event Method Trigger Sentence
+ACE Attack STAR raid As the rebels embarked on a daring trek across the desert, they launched a surprise raid on the heavily guarded fortress, catching the enemy off guard.
+SNaRe shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
+SPEED Death STAR asphyxiation The hiker’s life was tragically cut short as asphyxiation occurred after she became stuck in the narrow cave crevice.
+SNaRe killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
+GENIA Binding STAR merge The regulatory protein’s ability to activate a specific region of the DNA triggers the merge of two proteins, leading to the modification of gene expression.
+SNaRe bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
+Table 7: Qualitative examples demonstrating STAR and SNaRe’s trigger and sentence generation quality.
+7.2 Comparison with Powerful LLMs
+To further highlight the efficacy of SNaRe, we compare it with zero-shot inference using more powerful recent LLMs for the ACE dataset in Table 4.
+As seen, Llama3-70B with SNaRe significantly outperforms stronger LLMs like GPT4o and thinking-based models like QwQ-32B by 8-10% Eve-I F1 and 8-12% Tri-C F1 scores, respectively.
+This highlights the strong efficacy of synthetic generation SNaRe over zero-shot inference even when using more powerful LLMs.
+7.3 Comparison with Data mixing
+Data-mixing Hoffmann et al.
+(2022); Xie et al.
+(2023) is a widely used technique to leverage complementary information across datasets to promote robust downstream model training.
+We mix data from our two baselines of Weak Sup and STAR as a hybrid baseline to compare with SNaRe.
+To keep the comparisons fair, we consider $N/2=25$ data instances per event type from each dataset.
+Results from Table 5 demonstrate how SNaRe beats the data-mixing based hybrid model by 5-6% F1, underlining the significance of our three-stage model design over simpler data-mixing.
+7.4 Analyzing domain drift and label quality
+ED models have a strong tendency to over-rely on lexical relations between triggers and events Tong et al.
+(2022).
+Thus, we compare the synthetic data triggers with the gold test triggers as a raw study of the domain drift of triggers in the synthesized data.
+Specifically, we extract triggers per event type in the synthetic datasets and measure the hit rate/overlap of the synthesized triggers with the gold set of triggers, as reported in Table 6.
+STAR’s low hit rate indicates the poor overlap with the gold triggers, which is a primary reason for its domain drift.
+Furthermore, the consistently stronger coverage of SNaRe explains its lower domain drift.
+Method Naturalness Event Annotation
+Relevance Quality
+STAR 3.1 3.4 3.1
+Weak Sup 4.2 - 2.9
+SNaRe 3.6 4.0 3.6
+Table 8: Human evaluation for sentence naturalness, relevance of event in generated sentence, and the annotation quality. 1 = worst, 5 = best.
+Method ACE SPEED GENIA
+EI TC EI TC EI TC
+SNaRe 57.4 50.2 44.6 31.5 35.2 28.9
+   + SFT LLM 55.2 51.7 46.9 35.8 36.7 29.1
+Table 9: Measuring performance improvement by fine-tuning an LLM on unlabeled train data for SNaRe.
+To further study the quality and relevance of the label, we performed a human evaluation.
+Specifically, a human expert in ED is tasked with scoring generations (between 1-5) on the naturalness of the sentence specific to the target domain, the relevance of the event to the semantic actions described in the generated sentence, and the annotation quality evaluating the selection of triggers (details in § D.3).
+We provide the averaged scores across the three datasets for 90 samples in Table 8.222Since Weak Sup annotates the unlabeled target domain sentences, event relevance is analogous to annotation quality and we do not explicitly evaluate it.
+Weak Sup has high sentence quality but poor label annotations; STAR suffers from poor event relevance indicating domain drift.
+Overall, SNaRe performs the best with high annotation quality and event relevance.
+7.5 Domain-adapted LLM Fine-tuning
+We fine-tune the base LLM for Narrator on the unlabeled target domain train data $D_{T}^{\prime}$ to better align the generated passages.
+Naturally, this can be applied only for smaller LLMs owing to fine-tuning costs.
+We present the results of fine-tuning Llama3-8B-Instruct on the unlabeled train data in Table 9.
+On average, we observe that target data fine-tuning additionally improves SNaRe by 0.5-2% F1.
+Qualitative studies indicate that the generated passages are distributionally closer to the target domain, further reducing domain drift.
+7.6 Qualitative analysis of generated data
+We provide qualitative evidence for SNaRe’s reduction in domain drift by Scout’s domain-specific triggers in Table 7 (more examples in Table 19).
+We compare with STAR, which uses LLM’s internal knowledge to generate the triggers.
+Lack of domain grounding often results in STAR’s triggers and sentences being misaligned (e.g.
+asphyxiation for death event related to pandemics) relative to the target domain.
+In contrast, SNaRe’s triggers are better aligned to the target domain corpus, resulting in better quality data and reduced domain drift.
 8 Conclusion and Future Work
-We introduce FIG, a hybrid forward-inverse generation approach for better domain-specific synthetic data in low-resource ED.
-Experiments on three diverse datasets reveal that forward generation suffers from noisy labels due to poor LLM reasoning, while inverse generation faces domain drift.
-FIG mitigates these issues, achieving superior domain alignment and cleaner data, leading to the best downstream ED performance in zero and few-shot settings.
-Future works can explore enhancing stronger domain alignment and extending FIG for multilingual generation.
-Acknowledgments
-We express our gratitude to Anh Mac, Po-Nien Kung, and Christina Chance for their valuable time, reviews of our work, and constructive feedback.
+We introduce SNaRe, a domain-aware synthetic data generation approach composed of Scout, Narrator, and Refiner.
+Utilizing Scout’s domain-specific triggers for synthesizing sentences, along with Narrator’s conditional generation and Refiner’s annotations, helps reduce domain drift and label noise.
+Experiments on three diverse datasets in zero-shot, few-shot, and multilingual settings demonstrate the efficacy of SNaRe, establishing SNaRe as a strong data generation framework.
 Limitations
@@
 There are other specialized domains for ED as well which can be explored as part of future work.
-Finally, our proposed method FIG makes a practical assumption of access to unlabeled data to procure target domain cues to guide the data generation.
+Finally, our proposed method SNaRe makes a practical assumption of access to unlabeled data to procure target domain cues to guide the data generation.
 However, for specific super-specialized domains or if data has privacy concerns, this may not be possible and our method may not be applicable here.
@@
 Ethical Considerations
-The theme of our work is to generate high-quality domain-specific data using Large Language Models (LLMs) using forward-inverse generations.
+The theme of our work is to generate high-quality domain-specific data using Large Language Models (LLMs).
 The inherent LLMs can have certain biases, which can lead to potentially harmful or biased generations.
-Furthermore, the LLM can introduce potential hallucinations in the annotations which can hurt the model performance.
+Furthermore, the LLM can introduce potential hallucinations in the annotations, which can hurt the model’s performance.
 We do not check or consider any bias/hallucination detection method as part of our work, as it is beyond the scope.
-So future users should take due consideration of this vulnerability.
-Our proposed method FIG utilizes unlabeled data as a basis to procure domain-specific cues.
-If there are any biases in this data, it can propogate to the downstream model as well.
-We provide a proof-of-concept about our method in this work but do not detect or rectify such biases.
-Since the inverse generation paradigm causes the LLM to generate sentences/passages, which can potentially be copied from the pre-training data the LLM has been trained on.
-This can potentially lead to copyright infringements and we do not consider any of such violations under consideration for our method.
+Future works should take due consideration of this vulnerability.
+Our proposed method SNaRe utilizes unlabeled data as a basis to procure domain-specific cues.
+If there are any biases in this data, it can propagate to the downstream model as well.
+We provide a proof-of-concept about our method in this work, but do not detect or rectify such biases.
+SNaRe’s Narrator utilizes LLMs to generate sentences/passages.
+However, as noticed in past work, LLMs can potentially copy these sentences from the pre-training data on which it has been trained.
+This can potentially lead to copyright infringements, and we do not consider any such violations under consideration for our method.
 Users should consider this vulnerability before usage in commercial applications.
 We would also like to mention and acknowledge that we have utilized AI chatbots to help with the writing of the work.
-Appendix A Additional details about FIG
-We provide illustrations about the various prompts used in FIG.
-For target structure extraction forward generation, we consider a two-prompt approach.
-The first prompt aims to identify if any events of interest are mentioned in the text as illustrated in Figure 6.
-It comprises the task definition, full event ontology with definitions, the task instructions, and the query sentence.
+Appendix A Additional details about SNaRe
+We provide illustrations of the various prompts used in SNaRe.
+For Scout, we consider a two-prompt approach.
+The first prompt aims to identify whether events of interest are mentioned in the passage, as illustrated in Figure 6.
+This prompt comprises the task statement, full event ontology with definitions, the task instructions, and the query sentence.
 The second prompt identifies the trigger corresponding to the event of interest, as illustrated in Figure 7.
 Here, we specify the task definition, the event ontology details, and the query text with the task instructions.
-To aid inverse generation for passage generation, we provide the task definition, the event ontology with event definitions, and the query comprising the sampled target structure.
+For Narrator, we prompt the LLM with the task definition, the event ontology with event definitions, and the query comprising the sampled target structure.
+The LLM is asked to generate a sentence using the specified trigger words while mentioning the events of interest.
 We illustrate this prompt in Figure 8.
-Finally, we provide the simplified one-prompt setup for forward generation utilized for data refinement in Figure 9.
-Appendix B Additional Experimental Setup Details
-B.1 Data Statistics
+Finally, we provide the Refiner prompt in Figure 9. This prompt aims to identify any missing event mentions from the passage.
+Appendix B Data Statistics
 We discuss details about our dataset in § 5.
@@
 (2013) as the external data for GENIA.
+Finally, we also provide some statistics about the multilingual splits of the ACE dataset utilized for the Arabic and Chinese zero-shot experiments.
+[^3] For Chinese, the average length indicates the average number of characters.
 We provide statistics about this data in Table 10.
@@
 GENIA2013 6,542 - 17.4
-Table 8: Data Statistics for the various test and unlabeled datasets used in our work.
+Multilingual Test Data
+ACE - Arabic 313 198 24.6
+ACE - Chinese 486 211 44.2
+Unlabeled Multlingual Train Data
+ACE - Arabic 3,218 - 26.1
+ACE - Chinese 6,301 - 45.5
+Table 10: Data Statistics for the various test and unlabeled datasets used in our work. # = Number of.
 Appendix C Implementation Details
-Here, we provide detailed implementation details for each component and models used in our work.
-We run most of our experiments on NVIDIA RTX A6000/A100 machines with support for 8 GPUs, while for GPT3.5, we make API calls through OpenAI.
+Here, we provide detailed implementation details for each component and the models used in our work.
+We run most of our experiments on NVIDIA RTX A6000/A100 machines with support for 8 GPUs, while for GPT3.5, we make API calls through OpenAI using Curator Marten et al.
+(2025).
 C.1 LLM-based Generation
-We provide details about the various hyperparameters for using LLMs in all the components of STAR and FIG.
+We provide details on the various hyperparameters for using LLMs in all the components of STAR and SNaRe.
 For Llama3-8B-Instruct and Llama3-70B-Instruct, we present the hyperparameters in Table 11; while Table 12 presents the hyperparameters for GPT3.5.
@@
 For LLM-based prompting, we simply add these examples in the prompt as in-context examples to help the model do better reasoning/generation.
-For inverse generation methods (STAR, FIG), we do not add the $k$ triggers to the extracted/generated trigger list, as it leads to a drop in model performance.
-This can be attributed to the presence of duplicate information as the trigger generation/extraction already accounts for the $k$ triggers.
-For passage generation/Weak Sup, we append the $k$ datapoints to the synthetically generated data to provide signals from the gold data.
+For STARand SNaRe, we do not add the $k$ triggers to the trigger list, as it led to a drop in model performance.
+This can be attributed to the presence of duplicate information, as the trigger generation/extraction already accounts for these gold triggers.
+We also append the $k$ datapoints to the synthetically generated data to provide signals from the gold data.
 C.3 Downstream Model Training
@@
 (2022) as our downstream model for evaluation, a generation-based prompting model that utilizes natural language templates.
-We implemented the DEGREE model under TextEE framework Huang et al.
+We implemented the DEGREE model under the TextEE framework Huang et al.
 (2024).
@@
 (2020).
-We choose the task of causal language modeling (i.e.
-continual pre-training) to perform domain adaptation on unlabeled in-domain sentences.
+We choose the task of causal language modeling (i.e., continual pre-training) to perform domain adaptation on unlabeled in-domain sentences.
 We utilize cross-entropy loss on the dev split of the unlabeled data to select the best model.
@@
 Table 16: Example passages of overly long and more stereotypical sentences generated when the domain is mentioned or references are added to the LLM prompt for STAR.
-Figure 5: Model performance for FIG as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
+Figure 5: Model performance for SNaRe as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
 D.1 STAR with domain-specific prompt
-A simple way to infuse domain specific information in inverse generation pipelines like STAR would be to add domain-related information in the prompts to the LLM.
+A simple way to infuse domain-specific information in past works like STAR would be to add domain-related information in the prompts to the LLM.
 We experiment with two such methods: (1) domain-mention, where we provide the target domain information in the prompt and ask the model to generate accordingly, and (2) domain-reference, where we use some examples from the unlabeled data in the prompt as reference sentences to better guide the passage generation.
@@
 As observed, the results are generally poor, with an average drop of 0.1-0.6% F1 for domain-mention and 3.1-3.8% F1 for domain-reference.
-This is majorly because LLMs over-compensate, producing longer and more stereotypical information in their generations which hurts the naturalness of the sentence and causes further domain drift.
+This is mainly because LLMs over-compensate, producing longer and more stereotypical information in their generations, which hurts the naturalness of the sentence and causes further domain drift.
 Furthermore, the LLM makes more errors in mentioning the event as a part of its reasoning, which is utilized to make the generation in the domain style.
 We provide some qualitative examples for such generations in Table 16.
-In some ways, it also puts into light and amplifies the gains obtained by doing target domain SFT for FIG as discussed in § 6.4.
-Figure 6: Prompt for stage 1 of forward generation for target trigger extraction.
-Figure 7: Prompt for stage 2 of forward generation for target trigger extraction.
-Figure 8: Prompt for inverse generation for passage generation.
-Figure 9: Prompt for forward generation for data refinement.
+In some ways, it also puts into light and amplifies the gains obtained by doing target domain SFT for SNaRe as discussed in § 7.5.
+Figure 6: Prompt for stage 1 of Scout.
+Figure 7: Prompt for stage 2 of Scout.
+Figure 8: Prompt for Narrator.
+Figure 9: Prompt for Refiner.
 D.2 Impact of different number of training samples
-We conduct a small analysis to study the impact of changing the number of generated samples on the downstream model performance for FIG.
+We perform a small analysis to study the impact of changing the number of generated samples on the downstream model performance for SNaRe.
 We present the results for Llama3-8B-Instruct in Figure 5.
-As observed, the performance continues to increase as we increase the data from $N=10$ to $N=100$ datapoints per event type.
-This promises that inverse generation will provide continued improvements by having larger control over data distribution.
+As observed, performance continues to increase as we increase the data from $N=10$ to $N=100$ datapoints per event type.
+This promises that data generation will provide continued improvements by practicing greater and better control over the data distribution.
 Figure 10: Illustration of the interface for the human evaluation of the synthetically generated data.
@@
 We majorly evaluate on three dimensions: (1) Sentence naturalness (SN): This metric judges whether the sentence seems grammatical, natural, and fits the domain of the target data.
-(2) Event Relevance (ER): this metric is computed only for inverse/hybrid generation methods.
+(2) Event Relevance (ER): This metric is computed only for generation methods that generate sentences from labels.
 This evaluation judges whether the sampled event and trigger are appropriately used to generate a sensible alignment with the target domain.
@@
 We illustrate the annotation interface in Figure 10 and provide some sample examples in Table 17.
+Method # Unlabeled Data Eve-I Tri-C
+Inference - 46.9 41.3
+STAR - 50.0 42.3
+Weak Sup 100% train 53.2 48.0
+Weak Sup 20% train 51.4 46.1
+Weak Sup 5% train 50.1 44.9
+SNaRe 100% train 58.1 53.8
+SNaRe 20% train 55.9 52.1
+SNaRe 5% train 55.1 51.2
+Table 18: Ablating the amount of unlabeled data utilized for the ACE dataset by the different data generation methods using Llama3-70B-Instruct and its impact on downstream model performance.
+D.4 Analysis in low-resource unlabeled data settings
+In rare cases of hyper-specialized domains or domains with privacy concerns, there is a possibility of having super-low availability of unlabeled data.
+We believe that our current experimentation already simulates this setting, where we use only 773 documents for GENIA and 1.6k tweets for SPEED.
+Secondly, by providing strong efficacy of external data relative to in-distribution data, we demonstrate how even mildly related unlabeled data can also work well for SNaRe.
+Finally, we provide a deeper study of varying the amount of unlabeled data for the ACE dataset for Llama3-70B-Instruct as an ablative study to further substantiate the efficacy of our method in Table 18.
+Lowering the number of unlabeled data reduces the model performance relative to using higher number of data samples.
+However, even with 5% unlabeled training data ($\sim$850 samples), SNaRe still outperforms the other baselines, showing the efficacy of our method.
 Dataset Event Method Trigger Sentence
@@
 slay The wealthy businessman filed a complaint against his former partner, leading to a court proceeding, just as the rival gang leader slays his arch-nemesis in a brutal battle.
-FIG shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
+SNaRe shooting As the rival businessman signed the contract, a sudden shooting erupted outside, causing chaos in the midst of the transaction.
 bombing As the city was rocked by a devastating bombing, thousands of protesters took to the streets to demand justice and peace.
@@
 drowning As she struggled to stay afloat, her heart stopped beating, and she succumbed to drowning in the icy waters.
-FIG killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
+SNaRe killed The patient’s feverish state was triggered when they tested positive for the virus, which ultimately led to their being killed by the rapidly spreading infection.
 lost As the pandemic spread rapidly across the globe, thousands of people lost their lives due to the deadly virus.
@@
 snap When the two proteins snap together, the binding of the complex inhibits the expression of the target gene by deactivating a specific region of the DNA molecule.
-FIG bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
+SNaRe bound During the phosphorylation of the enzyme, it bound to the DNA sequence, initiating the transcription process.
 translocation The protein translocation to the nucleus triggers the induction of gene expression.
 binds When the enzyme binds to the substrate, it activates the addition of a phosphate group to the target molecule, marking a crucial change in its function.
-Table 16: Comparison of generated triggers and sentences from STAR and FIG methods
-D.4 Additional Qualitative Examples
-In § 6.5, we discussed how FIG improves domain drift qualitatively relative to STARand provided some examples.
+Table 19: Comparison of generated triggers and sentences from STAR and SNaRe methods
+D.5 Additional Qualitative Examples
+In § 7.6, we discussed how SNaRe improves domain drift qualitatively relative to STARand provided some examples.
 Here, we provide more examples to further support that study in Table 19.
-This table further demonstrates how STAR can have a domain drift without a lack of domain-specific cues, while FIG is better here.
+This table further demonstrates how STAR can have a domain drift owing to a lack of domain-specific cues, while SNaRe is better grounded in the target domain.

### v3

@@
 Low-Resource Event Detection
-Tanmay Parekh      Yuxuan Dong      Lucas Bandarkar      Artin Kim I-Hung Hsu      Kai-Wei Chang      Nanyun Peng Computer Science Department, University of California, Los Angeles        Google {tparekh, violetpeng, kwchang}@cs.ucla.edu
+Tanmay Parekh Yuxuan Dong Lucas Bandarkar Artin Kim I-Hung Hsu Kai-Wei Chang Nanyun Peng Computer Science Department, University of California, Los Angeles Google {tparekh, kwchang, violetpeng}@cs.ucla.edu
 1 Introduction
@@
 2 Problem Definition
-We focus on the task of Event Detection (ED) Sundheim (1992); Doddington et al.
+We focus on the task of Event Detection Sundheim (1992); Doddington et al.
 (2004) for this work.
@@
 Some 3,000 people have been arrested since the disobedience campaigns began last week.
-[…]
+Justice: Arrest-JailConflict: Demonstrate
 In our work, we specifically focus on ED in diverse and specialized domains (e.g., biomedical), where procuring a training dataset $D_{T}$ of annotated data points is expensive, but unlabeled data $D_{T}^{\prime}$ is available.
@@
 First, Scout extracts and filters domain-specific triggers, then Narrator generates passages conditioned on these triggers.
-Finally, Refiner adds any missing annotations and sample $N$ data points per event for downstream training.
+Finally, Refiner adds any missing annotations and sample NN data points per event for downstream training.
 Data Generation for Information Extraction
@@
 (2023); Wang et al.
-(2023b) and transfer-learning Cai et al.
+(2023b); Parekh et al.
+(2025a) and transfer-learning Cai et al.
 (2024) has been explored, but their performance remains inferior to supervised models Huang et al.
@@
 This motivates efforts in LLM-powered synthetic data generation for low-resource ED.
+Although our work only demonstrates results on ED, we believe our work can be extended to other tasks outside information extraction as well, like question-answering Rajpurkar et al.
+(2016); Parekh et al.
+(2025b) and long-form generation Suvarna et al.
+(2024); Chen et al.
+(2020); Parekh et al.
+(2020).
 4 Methodology - SNaRe
-In this work, we focus on domain-aware synthetic data generation using Large Language Models (LLMs) to alleviate the need for expert-annotated training data.
+In this work, we focus on domain-aware synthetic data generation using LLMs to alleviate the need for expert-annotated training data.
 By generating a large dataset $D_{s}=\{(X,Y)\}$, we can train domain-specific downstream ED models using minimal supervised data.
@@
 Both label noise and domain drift hurt the synthetic data quality, in turn, diminishing the downstream supervised model performance.
-To mitigate these issues, we propose SNaRe [Uncaptioned image], a domain-aware high-quality data synthesizer, comprising three components: Scout, Narrator, and Refiner.
-Scout studies unlabeled target domain data $D_{T}^{\prime}$ to curate domain-specific triggers, in turn, reducing domain drift.
+To mitigate these issues, we propose SNaRe [Uncaptioned image], a domain-aware data synthesizer, that generate and verifies LLM generations to ensure high-quality data Hsu et al.
+(2024), comprising three components: Scout, Narrator, and Refiner: Scout studies unlabeled target domain data $D_{T}^{\prime}$ to curate domain-specific triggers, in turn, reducing domain drift.
 Narrator generates domain-specific sentences conditioned on Scout’s curated triggers, while Refiner adds additional annotations to ensure high-quality labels.
@@
 We provide our architectural diagram in Figure 2 and explain each component of our pipeline below.
+Figure 3: Prompt for stage 1 of Scout.
+Figure 4: Prompt for stage 2 of Scout.
+Figure 5: Prompt for Narrator.
+Figure 6: Prompt for Refiner.
 4.1 Scout
@@
 Specifically, trigger extraction involves a two-stage prompt setup: (1) The first stage is tasked with identifying and filtering possible event types mentioned in the target domain sentence, and (2) The second stage aims to find the most appropriate trigger word from the unlabeled sentence for each filtered event type.
-We illustrate this setup in § A.
+We provide the prompts for Scout in Figure 3 and  4.
 To ensure high-precision of the triggers, we develop an aggregation and filtering mechanism by incorporating corpus-level statistics.
@@
 Here, we simply prompt the LLM to find all event mentions in the generated sentence (illustrated in Figure 6) and append them to the original sampled $Y$.
-To avoid adding noisy labels, we only add annotations for newly discovered events.
+Since these new refined labels can be noisy, we avoid updating them for the already present events from the Scout, and only add them for newly discovered events.
 To further improve data quality, we apply an automated rule to remove passages that do not mention the target trigger.
@@
 All our experimental results are reported over an average of three runs.
-Additional details are provided in Appendix C.
+Additional details are provided in Appendix B.
 6 Results
@@
 SNaRe performs the best:
-On average, SNaRe beats STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points – demonstrating how domain-specific cues from unlabeled data aid in tackling domain drift.
+On average, SNaRe outperforms STAR by 17.3% Eve-I F1 and 16.3% Tri-C F1 points – demonstrating how domain-specific cues from unlabeled data aid in tackling domain drift.
 Compared to Weak Sup, SNaRe provides average gains of 3.6% Eve-I F1 and 3.3% Tri-C F1, suggesting how cleaner label quality can help improve model performance.
@@
 We present the Tri-C results for all the datasets for the Llama3-8B model in Figure 8.
-Similar to zero-shot results, we observe that SNaRe consistently beats all other baseline models.
+Similar to zero-shot results, we observe that SNaRe consistently outperforms all other baseline models.
 On average, SNaRe outperforms STAR and Weak Sup by 5.4% Tri-C F1 and 7% Tri-C F1 respectively.
@@
    – Scout 43.2 27.8 28.2
+   – Narrator 37.8 31.5 21.4
    – Refiner 47.4 23.3 22.0
-Table 3: Ablation study for SNaRe’s Scout and Refiner measured as Tri-C F1 performance across datasets.
+Table 3: Ablation study for SNaRe’s Scout, Narrator, and Refiner measured as Tri-C F1 performance across the three datasets.
 LLM + Method Eve-I Tri-C
@@
 For ablating Scout, we replace it by prompting LLM to directly generate triggers.
-We observe how both Scout and Refiner are critical, with average performance reductions of 3.8% F1 and 6% F1 upon removing the respective components.
+Ablating Narrator is similar to the Weak Supbaseline, while adding additional refiner and dataset statistics to remove noisy datapoints.
+We observe how all the components are critical, with average performance reductions of 3.8% F1, 6.6% F1, and 6% F1 upon removing the respective components of Scout, Narrator, and Refiner.
 Dataset Event Method Trigger Sentence
@@
 To keep the comparisons fair, we consider $N/2=25$ data instances per event type from each dataset.
-Results from Table 5 demonstrate how SNaRe beats the data-mixing based hybrid model by 5-6% F1, underlining the significance of our three-stage model design over simpler data-mixing.
+Results from Table 5 demonstrate how SNaRe outperforms the data-mixing based hybrid model by 5-6% F1, underlining the significance of our three-stage model design over simpler data-mixing.
 7.4 Analyzing domain drift and label quality
@@
 To further study the quality and relevance of the label, we performed a human evaluation.
-Specifically, a human expert in ED is tasked with scoring generations (between 1-5) on the naturalness of the sentence specific to the target domain, the relevance of the event to the semantic actions described in the generated sentence, and the annotation quality evaluating the selection of triggers (details in § D.3).
+Specifically, a human expert in ED is tasked with scoring generations (between 1-5) on the naturalness of the sentence specific to the target domain, the relevance of the event to the semantic actions described in the generated sentence, and the annotation quality evaluating the selection of triggers (details in § C.6).
 We provide the averaged scores across the three datasets for 90 samples in Table 8.222Since Weak Sup annotates the unlabeled target domain sentences, event relevance is analogous to annotation quality and we do not explicitly evaluate it.
@@
 Experiments on three diverse datasets in zero-shot, few-shot, and multilingual settings demonstrate the efficacy of SNaRe, establishing SNaRe as a strong data generation framework.
+Acknowledgments
+We express our gratitude to Anh Mac, Po-Nien Kung, and Christina Chance for their valuable time, reviews of our work, and constructive feedback.
+We thank the anonymous reviewers, area chairs, and the program committee for their reviews and feedback.
+This work was partially supported by the National Science Foundation CAREER award #2339766 and the Amazon AGI Research Award awarded to Nanyun Peng.
 Limitations
@@
 We would also like to mention and acknowledge that we have utilized AI chatbots to help with the writing of the work.
-Appendix A Additional details about SNaRe
-We provide illustrations of the various prompts used in SNaRe.
-For Scout, we consider a two-prompt approach.
-The first prompt aims to identify whether events of interest are mentioned in the passage, as illustrated in Figure 6.
-This prompt comprises the task statement, full event ontology with definitions, the task instructions, and the query sentence.
-The second prompt identifies the trigger corresponding to the event of interest, as illustrated in Figure 7.
-Here, we specify the task definition, the event ontology details, and the query text with the task instructions.
-For Narrator, we prompt the LLM with the task definition, the event ontology with event definitions, and the query comprising the sampled target structure.
-The LLM is asked to generate a sentence using the specified trigger words while mentioning the events of interest.
-We illustrate this prompt in Figure 8.
-Finally, we provide the Refiner prompt in Figure 9. This prompt aims to identify any missing event mentions from the passage.
-Appendix B Data Statistics
+Appendix A Data Statistics
 We discuss details about our dataset in § 5.
@@
 Table 10: Data Statistics for the various test and unlabeled datasets used in our work. # = Number of.
-Appendix C Implementation Details
+Appendix B Implementation Details
 Here, we provide detailed implementation details for each component and the models used in our work.
@@
 (2025).
-C.1 LLM-based Generation
+B.1 LLM-based Generation
 We provide details on the various hyperparameters for using LLMs in all the components of STAR and SNaRe.
@@
 Table 12: Hyperparameters for decoding using GPT3.5 model.
-C.2 Few-shot Implementation Details
+B.2 Few-shot Implementation Details
 For the few-shot setting, we can access additional $k$ datapoints per event type to aid better performance.
@@
 We also append the $k$ datapoints to the synthetically generated data to provide signals from the gold data.
-C.3 Downstream Model Training
+B.3 Downstream Model Training
 We choose DEGREE Hsu et al.
@@
 Table 13: Hyperparameters for DEGREE model.
-C.4 LLM Fine-tuning
+B.4 LLM Fine-tuning
 We discuss domain-adapted passage generation through LLM fine-tuning in § 7.5.
@@
 Table 14: Hyperparameters for LoRA fine-tuning Llama3-8B-Instruct.
-Appendix D Additional analyses
+Appendix C Additional analyses
 In this section, we provide additional analyses to support our main experiments.
@@
 EI: Event Identification F1, TC: Trigger Classification F1.
+Method Original ACE Synthetic ACE
+EI TC EI TC
+LLM Training 71.5 67.4 54.3 46.8
+Small Model Training 72.2 68.9 57.4 50.2
+Table 16: Measuring model performance improvement providing domain-specific cues in the form of domain-mention (mention) or domain sentence references (references) to the LLM for STAR.
+EI: Event Identification F1, TC: Trigger Classification F1.
+Number of ACE SPEED GENIA
+Triggers EI TC EI TC EI TC
+$t=5$ 55.9 48.9 43.2 29.9 32.2 26.5
+$t=8$ 57.9 49.7 45.1 30.8 35.2 28.9
+$t=10$ 57.4 50.2 44.6 31.5 34.8 28.1
+$t=12$ 56.1 49.3 44.3 31.2 33.9 27.4
+Table 17: Ablating the impact of the number of triggers for the Scout component on downstream model performance across the three datasets. EI: Event Identification F1, TC: Trigger Classification F1.
 ACE
@@
 Table 18: Example passages of overly long and more stereotypical sentences generated when the domain is mentioned or references are added to the LLM prompt for STAR.
-Figure 5: Model performance for SNaRe as keep change the number of generated datapoints $N$ using Llama3-8B-Instruct for the three datasets.
-D.1 STAR with domain-specific prompt
+Figure 9: Model performance for SNaRe as keep change the number of generated datapoints NN using Llama3-8B-Instruct for the three datasets.
+Trigger Selection EI TC
+Frequency Ranking (current) 57.4 50.2
+Sampling 54.7 47.3
+Weighted Sampling 55.2 48.1
+Reranking 49.6 37.9
+Minimum Count Filtering 52.4 44.8
+Clustering 48.0 40.7
+Table 19: Ablating downstream model performance on the trigger selection strategy for the Scout component on the ACE dataset. EI: Event Identification F1, TC: Trigger Classification F1.
+C.1 STAR with domain-specific prompt
 A simple way to infuse domain-specific information in past works like STAR would be to add domain-related information in the prompts to the LLM.
@@
 In some ways, it also puts into light and amplifies the gains obtained by doing target domain SFT for SNaRe as discussed in § 7.5.
-Figure 6: Prompt for stage 1 of Scout.
-Figure 7: Prompt for stage 2 of Scout.
-Figure 8: Prompt for Narrator.
-Figure 9: Prompt for Refiner.
-D.2 Impact of different number of training samples
+C.2 Impact of different number of training samples
 We perform a small analysis to study the impact of changing the number of generated samples on the downstream model performance for SNaRe.
@@
 This promises that data generation will provide continued improvements by practicing greater and better control over the data distribution.
+C.3 Training LLMs with synthetic data
+Since we generate synthetic data using LLMs, it would be natural to fine-tune LLMs to be better at Event Detection (ED).
+We conduct a small experiment to test and compare the model training for LLMs and small models.
+Specifically, we train a Llama3-8B-Instruct model using Low Rank Adaptation (LoRA) Hu et al.
+(2021) and compare it with the best performing small models like DEGREE Hsu et al.
+(2022) and TagPrime Hsu et al.
+(2023a).
+We conduct sets of training: (1) Original ACE training data, and (2) Synthetic data generated by SNaRe for the ACE domain.
+We summarize our results and findings in Table LABEL:tab:llm-training.
+Overall, we note how our synthetic data can be used to train LLMs and improve their performance iteratively as well.
+However, we would like to note how small language model still learn better and provide better scores than LLMs on this task.
+We would also like to note that our LLM training can be suboptimal, as we tried using the original hyperparameters and didn’t tune the hyperparameters much.
+We utilized Llama-Factory Zheng et al.
+(2024) for our LLM fine-tuning.
+C.4 Ablating the design choice for number of triggers
+We utilize $t=10$ for ACE and SPEED and $t=8$ for GENIA.
+Here, we provide additional experiments to justify this design choice.
+Specifically, we provide the performance of training smaller models using different number of triggers ($t$) for the Scout using Llama3-8B-Instruct LLM in Table 17.
+Overall, we observe that changing the number of triggers affects the performance slightly, and can be optimally chosen to improve performance.
+However, we note that even with different $t$, SNaRe still outperforms the other baselines.
+C.5 Ablating the design choice for trigger selection
+We utilize frequency-based trigger selection in the Scout, but this can lead to missing out on low-frequency, long-tail, rare triggers.
+Our major motive is to avoid the noisy triggers which would also have low frequency.
+In order to justify this design choice, we conduct experiments with various other trigger selection designs, specifically: (1) Sampling: Instead of filtering and selecting, we sample uniformly from the trigger list, (2) Weighted Sampling: We sample using the extraction frequency of the triggers as the weights, (3) Reranking: Since some triggers might not be extracted correctly, we rerank the triggers based on the number of occurrences in holdout set (basically avoiding the bias of extracting them using the LLM), (4) Minimum Count Filtering: Instead of filtering based on ranks, we simply remove all triggers with a set minimum count (as this long tail might be noisy) and sample from the remaining triggers, (5) Clustering: We use k-means clustering to form clusters of similar triggers and sample one trigger from each - thus, providing enhanced diversity.
+We compare these various methods for trigger selection with the Scout using Llama3-8B-Instruct LLM on the ACE dataset in Table 19.
+Overall, this study shows how the different methods, which introduce more noise and provide different ranges of the precision-recall balance, are worse in comparison to our existing frequency ranking method.
 Figure 10: Illustration of the interface for the human evaluation of the synthetically generated data.
@@
 Table 20: Illustration examples for the human evaluation metrics. SN: sentence naturalness, ER: event relevance, AQ: annotation quality.
-D.3 Human Evaluation Details
+C.6 Human Evaluation Details
 We conduct a small human evaluation to judge the quality of the synthetic data in § 7.4.
@@
 Table 21: Ablating the amount of unlabeled data utilized for the ACE dataset by the different data generation methods using Llama3-70B-Instruct and its impact on downstream model performance.
-D.4 Analysis in low-resource unlabeled data settings
+C.7 Analysis in low-resource unlabeled data settings
 In rare cases of hyper-specialized domains or domains with privacy concerns, there is a possibility of having super-low availability of unlabeled data.
@@
 However, even with 5% unlabeled training data ($\sim$850 samples), SNaRe still outperforms the other baselines, showing the efficacy of our method.
+Method Arabic Chinese
+EI TC EI TC
+Inference 33.0 18.7 25.0 21.0
+STAR 28.1 22.0 32.3 25.9
+Weak Sup 29.4 20.2 30.2 28.4
+SNaRe (ours) 38.4 31.9 39.6 29.0
+Table 22: Comparing SNaRe with zero-shot inference for Qwen3-8B for the ACE dataset.
+C.8 Additional zero-shot multilingual experiments
+In § 6.3, we demonstrated the utility of our framework for zero-shot multilingual setting using the Llama3-8B and Llama3-70B models.
+Here, we additionally show results using the Qwen3-8B LLM in Table 22.
+We note that SNaRe outperforms the other baselines for both languages and metrics, even with the Qwen3-8B LLM.
 Dataset Event Method Trigger Sentence
@@
 Table 23: Comparison of generated triggers and sentences from STAR and SNaRe methods
-D.5 Additional Qualitative Examples
+C.9 Additional Qualitative Examples
 In § 7.6, we discussed how SNaRe improves domain drift qualitatively relative to STARand provided some examples.